A Complete Bibliography of Applied Statistics: 2010–2019

Nelson H. F. Beebe Center for Scientific Computing

University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

26 June 2021 Version 1.26

Title word cross-reference

2 × 2 [19]. 3 [95]. 48 [95]. D [303, 288]. L [99].

-optimal [99, 303, 288].

AAA [438]. absence [355]. abstentions [237]. abuse [457]. accelerometer [445]. accident [394]. Accommodating [352]. account [185]. Accounting [172, 382, 463, 433, 489, 107, 237, 325]. accumulation [462]. accuracy [201, 183, 89]. acid [97, 454, 329]. acoustic [401]. acquired [44, 328, 413]. across [78]. activation [412]. activity [244, 472, 122, 231]. actuarial [480]. acute [130, 69]. adaptation [265]. adaptive [281, 156, 438, 291, 188, 1, 196, 98, 230, 302]. Adding [231]. Additive [472, 418, 227, 106, 448, 369]. adjustment [139, 194]. adjuvant [79]. Adolescent [311]. adolescents [199, 342]. advanced [62]. adverse [338]. aero [9]. aetiology [246]. Affinity [66]. after [211, 194]. Age [132]. Age-dependent [132]. Ageing [56]. agent [438, 294]. agents [175, 31, 436, 435, 428, 265]. ages [475]. aggregate [397]. Aging [235, 250]. Agreement [313]. air [278, 136, 143, 285, 362, 357, 369]. alcohol [411]. algorithm [111]. algorithms [101, 337]. allocations [384]. allogeneic [460]. allow [338]. alone [248]. along [83]. Alzheimer [300]. ambient [357]. America [334]. American [102]. analyses [4, 119]. Analysing [485, 238, 360]. Analysis [363, 311, 225, 150, 76, 229, 245, 171, 450, 15, 202, 124, 315, 181, 409, 26, 411, 129, 403, 33, 143, 244, 187, 365, 354, 397, 7, 146, 285, 328, 486, 211, 362, 75, 351, 154, 329, 215, 155, 324, 359, 210, 206, 401, 348, 480, 147, 264, 25, 318, 32, 427, 195, 72, 67, 228, 220, 446, 491, 86, 323, 476, 393, 342, 45, 194, 369]. analytic [116]. analytics [417]. angular [264]. animal [264]. animals [132]. anisotropy [469]. Antarctic [60]. antigenic [463]. antipsychotic [114]. antiretroviral [216]. aphids [17]. application [380, 171, 104, 300, 399, 410, 59, 149, 82, 379, 483, 46, 419, 358, 44, 453, 214, 9, 33, 391, 127, 83, 356, 143, 290, 472, 311, 17, 350, 365, 354, 144, 148, 334, 462, 328, 225, 50, 477, 37, 140, 154, 296, 312, 215, 150, 484, 289, 387, 459, 266, 336, 309, 448, 261, 88, 299, 318, 5, 482, 77, 195, 76, 471, 234, 389, 164, 170, 6, 342, 320, 198, 458, 369]. applications [326, 481, 465, 64, 221, 417, 427, 263, 35, 413, 308]. applied [49, 333]. approach [386, 301, 136, 21, 253, 378, 192, 374, 398, 187, 190, 267, 48, 421, 138, 352, 494, 497, 106, 178, 336, 346, 158, 375, 28, 357, 364, 128, 353, 87, 228, 434, 249, 256]. approaches [293, 151]. appropriate [144]. Approximate [61]. approximation [426]. approximations [210, 58]. area [100, 388, 290, 37]. areal [205, 5, 10]. areally [238]. areas [91]. arguments [80]. arising [278, 341]. arms [27]. artefacts [339]. arthritis [80, 320, 370]. artificial [385]. aspects [79]. Assessing [252, 217, 89, 161, 236, 334]. Assessment [102, 327]. assignment [114]. assisted [433]. association [192, 55, 342]. associations [63, 136, 429]. assumption [436]. asthma [238]. Atlantic [482]. atomic [32]. attachment [57]. attribution [338]. atypical [185]. auctions [43]. auditory [152]. Australia [236, 204]. Australian [56]. Author [30]. authors [51, 85, 125, 165, 208, 240, 282, 343, 414, 498]. auto [181, 467, 297, 156, 435, 238, 395, 446, 368, 212]. auto-correlated [212]. auto-regression [238]. auto-regressive [181, 467, 297, 156, 435, 395, 446, 368]. Automated [104]. Automatic [76]. auxiliary [139]. availability [135]. avalanche [231]. avoids [447].

B [480]. B-spline [480]. bands [49]. bandwidth [42]. banknotes [212]. barnacles [471]. based [142, 409, 408, 129, 159, 135, 143, 376, 439, 3, 286, 53, 71, 106, 163, 74, 207, 6, 139, 249, 119, 256, 230, 455]. baseline [323]. basis [442]. basket [382]. Bat [353]. batch [104]. batsmen [424]. Bayes [48, 473]. Bayesian [466, 82, 396, 175, 347, 326, 382, 463, 21, 488, 61, 314, 253, 388, 127, 290, 29, 17, 99, 398, 168, 187, 439, 226, 260, 418, 255, 12, 441, 102, 242, 227, 50, 211, 269, 55, 224, 351, 27, 154, 438, 329, 497, 304, 317, 484, 2, 210, 346, 464, 88, 79, 32, 303, 427, 195, 131, 72, 389, 204, 228, 446, 373, 476, 434, 20, 90, 367, 265, 460, 30, 98, 35, 198, 251, 54, 455]. beam [60]. behaviour [26, 184, 263, 40]. behavioural [48]. behaviours [244]. benefit [459]. Bertalanffy [475]. beta [298]. between [136, 429, 55, 401, 128]. between-biomarker [429]. Bias [140, 492, 129, 341, 225, 351, 42, 359, 489]. biclustering [35]. bike [449]. bilateral [313]. Binary [297, 399, 392, 422, 110, 437, 133, 296, 497, 433, 179, 257, 220]. biodiversity [353]. bioelectrical [377]. bioinformatics [151]. biologic [435]. biological [175, 99, 72]. Biomarker [329, 429, 293, 279, 271]. biomarkers [429, 491, 230]. biomechanical [496]. biomedical [4]. biosimilar [336]. birth [136, 234]. bivariate [110, 199, 219, 480, 470]. BLAST [382]. Blinded [115]. block [390, 364]. body [342]. bombs [137]. Boosting [377, 106]. bootstrap [398, 42]. borrow [336]. Borrowing [78]. Botswana [128]. Bradley [134, 65]. brain [409, 316, 93, 45]. breast [225, 79, 68, 198]. brick [62]. Britain [82]. Brownian [25]. burden [199, 255]. business [314, 388, 153]. busulfan [460].

calculation [454]. Calendarization [145]. calibrated [336]. Calibrating [453]. calibration [396, 419]. California [238]. call [353]. caller [455]. caloric [122]. cancer [481, 486, 138, 3, 225, 211, 442, 443, 2, 79, 207, 68, 428, 198]. candidate [55]. Capture [82, 327]. car [262, 406]. cards [324]. career [66]. Careful [447]. caries [305]. carriers [167]. Case [487, 182, 200, 222, 24, 23, 486, 53, 362, 31, 80, 490, 262, 58, 259, 246]. case-control [200, 246]. Case-only [487]. case-study [182, 80, 490, 262, 58, 259]. cases [432]. cash [324]. categorical [10]. categorization [182, 329]. Causal [248, 323, 292, 27, 324, 80, 69]. causality [326]. cause [290]. cause-specific [290]. causes [486]. cell [330, 442, 460]. cells [22]. censored [300, 97, 38, 167, 150, 305, 434, 413]. censoring [385]. census [91]. central [259]. chains [483]. challenges [151]. Champlain [462]. Change [102, 83, 190, 32, 325, 45]. change-point [32]. changes [253, 334, 91, 120, 112, 291, 335]. Chao [356]. characteristic [217, 37]. characteristics [235]. Characterizing [173, 495]. cheating [40]. chemotherapy [79]. childhood [246]. children [472, 77]. China [410]. choice [303]. chromosome [304]. chronic [362, 458]. chronological [173]. chute [41]. circular [152]. civil [268]. claims [298, 131]. class [124, 26, 110, 391, 246]. classes [133, 366]. classification [270, 64, 168]. classified [19]. Classifying [124]. cleft [83]. Climate [102, 419, 326, 334, 14]. climatic [231]. clinical [281, 175, 44, 487, 115, 254, 255, 328, 48, 224, 80, 336, 428, 431, 196, 271, 302, 114]. cluster [358, 48, 421]. Clustered [370, 273, 392, 472, 354]. clustering [450, 390, 316, 168, 223, 144, 146, 445]. co [390]. co-clustering [390]. cocaine [212]. coclustering [485]. cocoa [164]. coefficient [422, 258]. cognitive [108, 56, 248, 228]. Cohort [350, 222, 458]. coinfection [31]. collected [170]. collective [330]. colonization [221]. colour [355]. combination [133, 474, 434]. combinational [175]. combinations [29, 438, 294, 428, 30]. Combining [374, 102, 107, 222]. commensurate [397]. communities [335]. community [364]. comparative [193]. Comparing [465, 257]. comparison [124, 6, 57]. comparisons [351, 65]. compartment [203]. compatriot [272]. competing [477, 18, 459, 490, 176, 164]. complex [322, 314, 53, 312, 118, 367]. compliance [27]. complications [140, 490]. component [15, 244, 285, 152, 4, 369]. components [7, 120, 206, 393]. composite [375, 67]. compositional [486, 476]. compound [276]. computation [61]. computational [191]. computationally [478]. Computer [419, 71, 266, 452]. conditional [467, 267, 156, 435, 213, 238, 373]. conditions [396, 41, 179]. condom [253]. confidence [49, 274, 456]. configurations [233]. confirmatory [103, 302]. confounding [292, 280, 350, 468]. Congressional [116]. connected [309]. consecutive [234]. Constrained [111]. Constructing [408]. consumption [390, 135]. contact [158]. contacts [13]. Contents [166, 209, 241, 283, 344, 415, 499]. context [297]. Contextual [366]. Continuous [25, 437, 481, 115, 23, 433, 65, 428]. contrast [203, 45]. contrast-enhanced [203]. contributions [231]. control [200, 222, 453, 252, 23, 201, 384, 474, 53, 246]. controlled [24, 247]. convolution [22]. cooled [62]. copula [423, 29, 421, 18, 317, 262, 373, 30]. copy [269]. core [64, 122]. correct [183]. correction [140, 359]. Correlated [387, 399, 351, 305, 220, 478, 212]. correlation [363, 318, 470]. cosmological [42]. cost [12, 216]. cost-effectiveness [12]. count [365, 184, 50, 448, 263]. Counting [274, 320]. counts [467, 187, 354]. county [360]. coupled [96, 482, 157]. course [124]. covariance [275, 172]. covariances [120]. covariate [300, 23, 362, 103, 194]. covariate-specific [23]. covariates [182, 481, 117, 147, 407, 179, 263, 370, 413]. Cox [468, 427, 447, 245, 323]. cracking [62]. crash [262]. credit [478]. Creutzfeldt [46]. crime [326, 407]. criminal [184]. criminology [28]. critical [298, 121, 471]. criticism [452]. Crohn [490]. crop [480]. Cross [289]. Cross-validatory [289]. cure [3, 232, 176]. cured [38]. current [477, 375]. curve [49, 37]. curves [49, 390, 465, 121, 146, 471]. cutting [459]. cycle [347, 190, 153]. cycles [437]. cylindrical [375]. cytostatic [265]. cytotoxic [428].

damage [276, 320]. Data [466, 357, 380, 363, 171, 19, 313, 182, 200, 273, 423, 371, 191, 34, 411, 483, 392, 84, 326, 214, 253, 143, 199, 290, 244, 472, 311, 376, 13, 398, 97, 187, 365, 223, 354, 226, 397, 340, 285, 310, 486, 81, 50, 211, 477, 38, 352, 445, 75, 167, 497, 106, 132, 215, 152, 150, 359, 484, 118, 205, 39, 206, 218, 402, 305, 336, 346, 448, 401, 348, 261, 158, 238, 264, 25, 5, 77, 195, 470, 262, 232, 58, 485, 179, 65, 228, 10, 446, 47, 373, 170, 406, 491, 476, 434, 368, 212, 393, 73]. data [369]. Datalink [214]. death [486, 352, 360, 176]. deaths [486]. debit [324]. decision [408, 87]. decline [56, 248]. Decomposing [267]. decomposition [459]. decrease [324]. default [242]. Defibrillator [76]. deficiency [44, 328, 413]. degraded [84]. Delaunay [100]. delay [298]. delayed [341, 465, 196]. delays [312]. demand [449, 324]. dementia [89]. dendrite [202]. dengue [467]. density [60, 495, 88]. dental [365]. Dependence [262, 24, 452, 259]. dependent [350, 18, 132, 73]. deposition [5]. derived [2]. descriptors [301]. Design [288, 182, 281, 175, 382, 99, 105, 254, 439, 255, 451, 219, 454, 436, 433, 435, 309, 147, 67, 428, 440, 431, 98, 271, 73, 40]. designs [59, 287, 488, 338, 111, 224, 438, 303, 95, 230, 302]. detect [304, 54]. Detecting [423, 341, 469, 452, 405, 112]. detection [421, 329, 377, 335, 256, 457]. determination [48]. determine [437]. Determining [137]. developed [19]. development [138]. device [135]. diagnoses [395]. diagnosis [211, 89]. diagnostic [222, 193, 327, 257]. diagnostics [266]. dichotomous [487]. diet [197]. dieting [247]. differences [205, 401]. different [231, 41]. differential [376, 187, 464, 68]. differentially [310, 346, 220]. diffusion [261]. digit [332]. dimensional [200, 315, 386, 481, 106, 484, 88, 493, 251]. direct [247, 40]. directed [488]. directional [172]. Dirichlet [81, 417]. Disaggregated [10]. discontinuation [39]. discontinuous [65]. Discovering [391]. discovery [6]. discrepancy [87]. Discrete [448, 250]. disease [410, 191, 46, 13, 189, 384, 167, 317, 387, 360, 291, 490, 157, 458, 369]. diseases [174]. disorder [100, 411]. disorders [171]. disparities [16]. dispersion [100]. distance [60]. Distributed [420, 218]. distribution [82, 298, 167, 163, 74, 261, 164]. distribution-free [261]. distributional [227, 112]. distributions [113, 1]. disturbing [26]. diversity [335]. DNA [159, 297, 186, 47, 54]. Do [324]. does [436]. Domain [348]. Domain-selective [348]. dominated [111]. Dose [441, 175, 432, 488, 437, 29, 133, 254, 474, 443, 438, 258, 294, 428, 440, 265, 30]. dose-response [258]. dose-schedule [254]. doses [442]. dosing [460]. double [182]. downscaling [204]. drifter [243]. driven [489]. drivers [148]. driving [330]. dropout [280, 248, 352, 180]. drought [378]. drug [29, 474, 213, 30]. drug-induced [213]. drugs [489]. dual [438, 40]. dual-agent [438]. dunlin [197]. Dupuytren [317]. durations [459]. dye [54]. dye-labelled [54]. Dynamic [134, 367, 153, 371, 44, 279, 206, 485, 229, 114, 179, 203, 393, 370, 458]. dynamic/transition [179]. dynamics [70, 316, 374, 189, 43, 407].

early [293, 441, 443, 431, 247]. earthquake [218]. eastern [236]. echo [60]. echo-sounder [60]. echolocation [353]. Ecological [339]. ecology [376, 264]. economic [142, 144, 87]. edge [489]. Editors [126]. education [128]. effect [313, 15, 272, 34, 292, 408, 391, 122, 350, 248, 459, 131, 69]. effectiveness [12, 114]. effects [149, 487, 127, 465, 252, 81, 27, 239, 258, 172, 464, 103, 76, 176, 68, 130, 491, 476, 263, 370, 161, 247, 57]. efficacy [435, 90]. efficiency [191]. efficient [478]. egocentric [158]. electricity [390]. electrocardiograph [146]. electroencephalography [484]. emerging [384]. empirical [82, 74]. Emulating [189]. emulation [322, 496, 266]. end [292, 487, 439, 150, 433]. endogenous [179]. engine [9]. enhanced [203]. Ensemble [490, 21, 188]. environmental [96, 50, 112, 41, 389]. epidemic [46]. epidemiological [322]. epidemiology [278, 50, 80, 28, 389]. epistatic [423]. equality [121]. equation [394, 376, 464]. equations [214]. error [278, 475, 150, 210, 183, 139]. errors [141, 107]. erythematosus [387]. estimate [253, 158]. estimated [492, 36, 453]. estimates [36, 385, 74, 139]. Estimating [371, 396, 483, 292, 60, 316, 199, 13, 350, 3, 42, 27, 167, 114, 128, 109, 69, 93, 247, 40, 149, 136, 214, 78, 43, 248, 468, 291]. Estimation [475, 22, 306, 142, 82, 386, 394, 388, 337, 310, 242, 286, 71, 312, 150, 2, 88, 79, 357, 164, 174, 373]. estimator [323]. estimators [356, 77]. Euphausia [60]. evaluate [216]. Evaluating [47, 130, 23, 258, 77]. Evaluation [159, 143, 74, 103, 58, 212]. event [136, 442, 490, 288, 470, 491, 393]. event-related [288]. events [24, 338, 234]. evidence [114, 47, 212]. evolutionary [301]. evolving [57]. exact [99]. examining [205]. exciting [407]. existence [320]. expectile [128]. experimental [463, 219]. experiments [59, 70, 99, 105, 451, 111, 178, 309, 288]. expert [231]. Exploratory [160, 63]. Exploring [449, 320, 487, 401, 68]. exponential [310]. exposure [8, 149, 362, 359, 88]. exposures [278, 77, 357]. expressed [346]. extension [118]. extent [79]. external [422]. extraction [104]. extrapolation [359]. extremal [259]. extreme [396, 334, 462, 385, 289, 213].

facilitate [337]. facilities [476]. factor [171, 116, 2, 377]. factor-specific [377]. factorial [182]. factors [456]. failure [225]. failures [219]. fair [480]. familial [186]. families [138]. family [310]. farming [325]. Fast [496, 119]. feature [104]. fecundity [215]. feeding [247]. fermentation [104]. fertility [448, 128]. field [6]. fields [33, 375]. filter [94]. financial [315, 160]. find [144]. finding [175, 432, 488, 29, 133, 254, 441, 474, 438, 440, 265, 30]. findings [81]. fingerprint [469]. Finite [155, 483, 97, 3]. fish [270]. fixation [355]. flashes [235]. Flexible [169, 218, 86, 43, 106, 258]. flood [462]. flows [141, 41]. follicle [235]. foot [410]. force [32]. Forecasting [486, 141, 21, 113]. forecasts [290, 188]. forests [487]. formation [229]. Forward [256]. fraction [3, 176]. fractures [354]. Frailty [57, 174]. framework [63, 143, 14, 454, 27]. France [259]. fraud [457]. free [439, 255, 2, 261, 434, 93]. French [46]. frequency [121]. frost [236]. fruits [164]. Fully [356, 154]. function [252, 219, 71, 74, 65]. function-on-scalar [252]. Functional [445, 4, 393, 15, 390, 396, 408, 108, 223, 146, 111, 402, 348, 261, 377, 288, 446]. functioning [64]. functions [9, 121, 362, 167]. Fusing [5]. fusion [268]. fuzzy [374].

gamma [163]. gaps [267]. gas [62]. gas-cooled [62]. Gaussian [34, 226, 317, 427, 447, 493, 245, 412]. gBOIN [433]. gene [37, 72]. General [133, 356, 264, 313]. Generalized [106, 369, 59, 275, 467, 214, 298, 192, 17, 418, 163, 448, 113, 65]. Generalizing [178]. generation [59]. generational [366]. genes [55]. genetic [192, 260, 111, 318, 86, 456]. genetically [138]. genomewide [454]. genomic [6]. genotype [63, 423, 487]. genotype-specific [487]. geoadditive [20]. geodesics [83]. geographical [82, 301]. geometric [297]. Geostatistical [11]. geostatistics [425]. Gibbs [405]. girl [247]. glacial [347]. global [301, 357]. goals [195]. grades [433]. graph [310]. graphene [277]. graphical [423, 317, 412]. graphics [160]. graphite [62]. gravity [189, 249]. gravity-based [249]. Great [82]. Group [101, 214, 233, 255]. grouped [359]. grouping [339]. groups [81]. growth [475, 17, 7, 469, 2, 164, 277]. guidance [256].

H1N1 [463]. halving [101]. hand [410]. Handling [179]. harmonic [59]. hazard [481, 323]. hazards [468, 232, 323]. Health [476, 181, 149, 362, 87, 249, 263, 311]. healthcare [74]. healthy [45]. heart [360]. heat [349]. heatwaves [259]. heritability [386]. herpes [31]. heterogeneities [174]. heterogeneity [380, 382, 391, 269, 325, 179, 161, 57]. heterogeneous [101, 442]. Heteroscedastic [238, 467]. heteroscedasticity [141]. heuristic [337]. Hidden [321, 82, 46, 495, 226, 304, 375, 157]. Hierarchical [162, 233, 149, 463, 21, 290, 252, 260, 190, 102, 154, 329, 2, 131, 357, 204, 249, 198]. hierarchically [493]. hierarchies [119]. high [200, 315, 386, 138, 137, 106, 484, 88, 493, 335, 251]. high-risk [138, 137]. higher [452]. historical [258, 336, 377]. histories [362]. History [322, 305]. HIV [350]. homogeneous [226, 367]. hormone [235]. hospital [168, 211]. hospitalization [238]. hot [235]. household [158]. Houston [349]. human [322, 44, 253, 97, 216, 31, 88, 395, 413]. humanitarian [266]. hurricane [276]. hybrid [222].

identification [159, 463, 438, 353, 57].

identify [244]. Identifying

[293, 39, 346, 72]. ignorability [365].

ignorable [492, 280, 117, 244, 248, 75, 180]. II [254, 439, 137, 436, 435, 294, 428, 98]. illness [298]. image [409, 33, 45]. image-based [409]. images [4]. imaging [111, 261, 288, 203, 446, 90]. immune [44, 328, 413]. immunodeficiency [322, 44, 253, 97, 216, 31, 395, 413]. immunotherapy [465, 439]. imperfectly [230]. Implementing [313]. implications [485]. importance [200, 456]. improve [385, 2]. Improved [335]. Improving [386, 463]. in-hospital [168]. incautious [447]. incidence [91]. income [170]. incomplete [358]. incompletely [19]. inconstancy [194]. incorporate [301]. incorporated [340]. indentation [32]. independent [64]. Index [51, 85, 125, 165, 208, 240, 282, 343, 414, 498, 342]. indicators [142, 183]. Indices [492, 378, 143]. indirect [61]. individual [24, 397, 67, 161, 57]. individualization [441]. induced [213, 45]. induction [130]. inequalities [481]. infection [380, 92, 22]. infectious [410, 191, 13, 189, 384, 174]. infer [189]. Inference [197, 1, 466, 191, 496, 11, 61, 330, 17, 105, 426, 248, 269, 324, 402, 464, 299, 364, 447, 478]. Inferences [207]. inferencing [367]. inferiority [194]. inferring [260]. inflated [306, 365, 239, 263]. inflation [354, 494]. influences [231]. influenza [463]. Information [331, 284, 295, 307, 319, 345, 361, 372, 383, 400, 416, 430, 444, 461, 479, 281, 422, 78, 436, 107, 336]. Information-adaptive [281]. informative [371, 358, 352]. Informing [422]. initial [179]. initialization [94]. initiating [216]. inject [489]. injury [394, 213, 69]. input [445]. instrumental [468, 128]. insurance [298, 480, 131, 406]. integer [326, 467]. integer-valued [326, 467]. integrated [94, 426, 210, 58]. Integration [497, 357]. integrative [466]. intelligence [314]. intensities [370]. Interaction [68, 420]. interactions [260, 405, 318, 157]. interglacial [347]. intermittently [482]. international [64]. interpolating [145]. interpolation [333]. interpreting [14]. Interval [38, 224, 150, 305, 164]. Interval-censored [38, 150, 305]. Intervention [76, 114, 68]. intratumour [269]. intravenous [460]. inverse [251]. inversion [198]. Investigating [180, 211, 120]. involving [27]. ionospheric [121]. isotonic [127]. isotope [197]. Issue [331, 284, 295, 307, 319, 345, 361, 372, 383, 400, 416, 430, 444, 461, 479, 177, 308, 431]. issues [52, 454]. Italian [364]. item [184].

Jakob [46]. Janeiro [333]. Joint [437, 56, 477, 215, 342, 250, 340, 352, 130, 90, 320]. jump [256].

Kalman [94]. Kaplan [465]. key [151]. killer [442]. kind [298]. kinetic [99]. kinetics [277]. kinship [159]. know [274]. krill [60].

labelled [54]. labour [477]. lag [420]. Lagrangian [243]. lags [218]. Lake [462]. Landmark [458]. languages [401]. Laplace [426, 210, 58]. large [301, 84, 419, 299, 41]. large-scale [301, 299]. lasso [441, 453]. late [122]. Latent [26, 185, 450, 124, 181, 390, 411, 382, 110, 226, 148, 81, 229, 493, 246]. left [496]. length [190]. leukaemia [130]. level [397, 360, 5, 95, 370]. levels [97, 334]. licensed [148]. life [124, 279, 305, 485]. lifetime [171, 164]. light [366]. likelihood [129, 394, 192, 337, 267, 75, 158, 375, 65]. likelihood-based [129]. limited [253]. line [43, 384]. Linear [300, 182, 59, 8, 275, 386, 92, 451, 296, 299, 413, 73, 458]. Linguistic [15]. linked [119]. lip [83]. liver [213]. living [248]. load [49]. local [484]. localization [152]. Locally [33, 156]. location [379, 106]. locations [355]. Log [296, 427, 447, 245]. log-Gaussian [427, 447, 245]. Log-mean [296]. logistic [392, 201, 140]. logit [337, 113]. logit-normal [113]. long [44, 235]. long-term [44, 235]. Longitudinal [181, 311, 108, 56, 45, 26, 411, 97, 248, 473, 279, 352, 215, 206, 162, 86, 393, 263, 342, 458, 250]. lottery [272]. low [69]. lower [399]. lung [207, 69]. lupus [91, 387].

macroeconomic [371]. magnetic

[111, 288, 203, 446, 90]. management [471]. Managing [87]. mapping [349]. Marginal [392, 354, 232, 342]. marginalized [239]. marked [132]. marker [230, 271]. marker-adaptive [230]. marker-stratified [271]. markers [23]. Markov [450, 483, 46, 467, 495, 226, 355, 304, 221, 375, 157, 321, 6]. mass [34, 42, 155, 342]. match [424]. matched [19, 200, 23]. matching [322, 313]. material [396, 32]. Mathematical [31]. mathematics [391]. Mauritius [489]. maximum [337]. meal [299]. mean [121, 296, 470]. means [82]. measles [92]. measured [475, 437, 215, 147, 230]. measurement [141, 278, 210, 183]. measurements [108, 38, 279, 257]. measures [152, 4]. Measuring [237]. Mechanism [340]. mechanisms [330]. median [36]. Mediation [365, 323]. medical [457]. Medicare [16]. Meier [465]. membership [76]. memory [79, 427]. menopause [190]. menstrual [190]. mental [171]. meta [129, 403, 397, 351, 154, 329, 427]. meta-analysis [129, 403, 397, 351, 154, 329, 427]. metabolomics [206]. metaregression [389]. method [214, 328, 184, 103, 480]. methodological [52]. methodology [137, 431]. Methods [425, 432, 60, 23, 53, 39, 198, 54]. Mexican [199]. mice [122]. microbial [335]. microelectronics [127]. microscopy [32]. misaligned [285, 402]. misclassification [271]. misclassified [220, 491]. mismeasured [362, 413]. missing [244, 311, 398, 340, 75, 76]. Mixed [34, 258, 15, 182, 386, 411, 144, 473, 351, 464, 299, 478]. mixed-type [144]. mixing [197]. Mixture [63, 171, 181, 358, 199, 3, 167, 304, 132, 155, 2, 205, 1]. mixtures [466, 34, 483, 97, 417, 303, 67, 493, 47]. mobility [421, 180]. model [171, 322, 236, 390, 396, 347, 419, 222, 422, 333, 496, 463, 475, 495, 110, 481, 143, 199, 290, 376, 306, 116, 268, 334, 189, 43, 381, 102, 14, 140, 18, 239, 304, 327, 152, 258, 2, 433, 205, 206, 266, 305, 346, 448, 360, 264, 291, 131, 232, 485, 176, 357, 157, 203, 204, 277, 65, 478, 323, 434, 90, 265, 263, 35, 413, 73, 458]. model-assisted [433]. model-based [143]. model-free [434]. Modelling [380, 399, 332, 44, 97, 190, 462, 235, 234, 259, 142, 63, 182, 410, 94, 8, 34, 134, 84, 135, 314, 33, 437, 297, 56, 354, 226, 260, 268, 96, 340, 138, 184, 231, 477, 91, 352, 154, 62, 155, 317, 484, 218, 213, 407, 28, 262, 364, 10, 321, 249, 370, 457, 251, 153]. models [466, 450, 15, 273, 181, 423, 371, 59, 149, 275, 386, 379, 191, 411, 46, 358, 44, 92, 326, 420, 467, 197, 24, 22, 133, 449, 17, 99, 398, 426, 168, 226, 418, 337, 190, 451, 12, 148, 310, 355, 473, 81, 3, 156, 286, 55, 296, 329, 31, 468, 106, 132, 215, 221, 317, 185, 435, 210, 169, 80, 387, 162, 464, 103, 145, 238, 405, 375, 299, 25, 377, 482, 262, 66, 229, 395, 179, 87, 243, 174, 170, 6, 20, 368, 246, 320, 370, 413, 369, 412, 250]. Moderated [65]. moderators [408]. modified [111]. molecular [198]. molecularly [436]. monetary [256]. monitoring [115, 348, 353]. monotonicity [436]. morphological [146]. mortality [290, 418, 362, 207, 16]. motion [25]. motor [252]. mouth [410]. movement [330, 376, 264]. mover [320, 370]. Movers [325]. moving [393]. Multi [481, 60]. multi-beam [60]. Multi-dimensional [481]. Multicenter [76]. Multiclass [368]. multicountry [272]. multilevel [299, 170]. multimodal [376]. multimorbidity [296]. multinomial [354]. multiobjective [111]. multiphase [89]. multiple [149, 192, 133, 442, 349, 387, 103, 261, 77, 114]. multiple-assignment [114]. multiplexed [54]. Multipollutant [278]. multirater [195]. multiscale [268]. multispecies [381, 245]. multistate [138, 221, 80, 387, 305, 370]. multistore [368]. multitreatment [27]. Multitype [202]. Multivariate [141, 8, 275, 411, 426, 146, 266, 171, 403, 378, 285, 227, 225, 55, 205, 402, 360, 405, 389, 170, 245]. mutation [167, 455]. mutations [269].

nanoparticle [100]. national [272, 311]. nations [450]. natural [442]. negative [78]. neonates [440]. nested [363, 426, 215, 210, 58]. Network [6, 313, 202, 315, 223, 397, 310, 452, 308, 367]. Network-based [6]. networked [318]. networks [314, 311, 421, 312, 169, 309, 162, 158, 229, 57]. neuroimaging [427]. neuron [202]. neuronal [108]. nevi [77]. newly [148]. Non [273, 473, 474, 389, 139, 492, 59, 8, 419, 92, 453, 333, 280, 244, 226, 451, 248, 111, 55, 27, 468, 489, 482, 195, 471, 373, 321, 180, 367, 460, 194, 413]. non-compliance [27]. non-dominated [111]. non-homogeneous [226, 367]. non-ignorable [492, 280, 244, 248, 180]. non-inferiority [194]. non-linear [59, 8, 92, 451, 413]. Non-parametric [273, 473, 474, 389, 55, 195, 471, 373, 460]. non-probability [492, 453]. non-proportional [468]. Non-response [139]. non-stationary [419, 333, 482]. non-uniform [489]. non-volcanic [321]. normal [113]. North [334, 102, 482]. novel [378, 285]. nuisance [115]. number [269]. numbers [274]. nurse [362].

Objective [242]. objects [7]. observation [221, 370]. observational [292, 214]. observations [60, 164, 321]. observed [423]. ocean [495, 289, 243]. ocular [355]. odds [140]. offenders [173]. oligonucleotides [54]. on-line [43, 384]. oncology [175, 29, 30]. one [306]. one-inflated [306]. only [84, 487]. onset [122, 167]. operating [37]. Optimal [59, 233, 287, 193, 451, 219, 384, 221, 309, 230, 437, 99, 293, 224, 438, 303, 288, 114]. Optimization [135, 337, 71]. Optimizing [442, 294, 460]. Optimum [105]. order [328, 417]. ordered [273]. ordering [213, 207]. orders [446]. Ordinal [148, 340, 195, 485, 294]. ordinally [64]. oscillation [482]. osteoporotic [354]. outcome [437, 4, 73]. outcome-dependent [73]. outcomes [422, 403, 108, 248, 351, 490, 220, 294, 161, 250]. outputs [396, 419, 102, 14]. Ovarian [235]. overall [130]. overdose [474]. overexpressed [304]. overhaul [9].

paediatric [93]. PairClone [455]. paired [199, 65]. pairs [19, 455]. pancreatic [138, 211]. panel [170, 180]. parameter [496, 337, 12, 42, 215]. parameters [115, 13]. Parametric [163, 294, 273, 473, 474, 55, 195, 471, 389, 373, 460]. Pareto [142]. Parliament [364]. partial [318, 207]. Partially [246, 423, 73]. participation [249]. passive [366]. pathway [412]. pathways [72]. patient [382, 397]. patients [26, 442, 69, 93, 161]. Pattern [358, 84]. patterns [332, 449, 116, 349, 452, 405, 245]. penalized [278, 394, 481, 108, 158, 480, 364, 65]. Penn [235]. people [489]. Peptide [404]. Percentile [74]. Percentile-based [74]. performance [74]. periodontal [273]. persistent [26]. pertussis [199]. Phase [435, 98, 254, 439, 441, 443, 224, 436, 433, 294, 428, 431, 434, 338]. phonetic [401]. phosphorylation [2]. physical [244, 472, 122]. pitch [15]. plantation [191]. plants [7]. platelet [2]. platelet-derived [2]. platform [443]. plot [67]. pneumococcal [221]. Point [84, 202, 135, 449, 439, 190, 137, 349, 405, 407, 28, 32, 5, 447, 245]. points [292, 487, 150, 433, 471]. Poisson [306, 239, 276]. policy [256]. political [453]. polling [453]. pollutants [149, 420]. pollution [278, 136, 285, 362, 357, 369]. population [94, 356, 143, 17, 306, 3, 53, 320, 16]. population-based [3, 53]. populations [101, 78, 55]. portfolio [71]. positive [306, 78]. post [391, 286, 188]. post-processing [286, 188]. post-treatment [391]. posterior [426]. potential [376, 161]. Power [397, 135, 454, 336, 113, 406]. practical [52]. practices [247]. prealigned [7]. precipitation [334]. precision [460]. predictability [256]. predicted [117]. Predicting [216, 41, 249, 90, 278, 234]. Prediction [298, 493, 49, 422, 279, 497, 490, 395, 393, 458]. predictive [78, 201, 406, 98]. predictiveness [23]. predictors [408, 235]. Preface [177, 308, 431]. preference [332]. preferential [11, 425, 57]. pregnancies [234]. pregnancy [358]. preliminary [77]. prescription [39, 457]. presence [141, 84, 280, 244, 168, 355, 225, 468, 271]. presence-only [84]. present [339]. preterm [136, 234]. prevalence [253, 40]. prevention [216, 195]. price [43]. pricing [406]. principal [15, 244, 7, 285, 324, 120, 206, 4, 393, 369]. prior [336, 447]. prioritization [443]. prioritizing [195]. priors [397, 484, 207]. private [310]. Probabilistic [244, 186, 206, 113, 353]. probabilities [489, 109]. probability [492, 453, 242, 98]. probe [198]. probit [116, 478]. problems [314, 133, 251]. procedure [468, 89]. process [104, 202, 70, 135, 449, 81, 137, 276, 417, 427, 320]. processes [380, 301, 312, 387, 238, 112, 407, 447, 245]. processing [286, 188]. profile [348]. profiles [104, 26, 159, 244]. Program [102]. programme [68, 249]. progression [439, 255, 2, 387, 93]. progression-free [439, 255, 2, 93]. Projecting [418]. propagation [312]. propensity [313, 178]. Properties [77, 396]. proportional [140, 468, 232, 323]. proportions [492, 354]. prostate [2]. protein [151]. proteomic [34]. providers [74]. proxies [107]. psoriatic [80, 320, 370]. public [312]. publication [129]. purity [123]. puzzle [494].

quality [143, 485, 357]. quality-of-life [485]. Quantifying [100, 176]. Quantile [279, 472, 267, 50, 71, 385, 163, 112, 342]. quantitative [55, 47, 90]. quantum [123]. Quasi [70, 192]. Quasi-experiments [70]. quasi-likelihood [192]. questioning [40]. questionnaire [77]. questions [184].

race [360]. Racial [16]. radiation

[127, 359, 45]. Radiologic [409]. rain [462]. rainfall [226, 204]. Raman [54]. random [487, 337, 310, 81, 239, 150, 147, 176, 6, 370]. random-effects [81]. random-field [6]. random-graph [310]. randomization [196, 98]. randomized [487, 48, 76, 114, 40]. randomly [300]. rank [118, 86]. ranked [229]. Ranking [456, 424, 366]. rate [150, 232]. rates [480, 360, 207]. ratio [36]. ratios [193]. reactors [62]. readmission [211]. Realtime [270]. recapture [82, 327]. receiver [37]. recently [19]. receptor [2]. recommendation [109]. reconciliation [52]. recording [341]. recovery [132, 79]. recruitment [281]. Rectangular [450]. recurrent [380, 24, 470]. Reduced [149, 489, 86]. reduction [440]. reference [207]. referenced [238]. refinement [404]. regime [495, 374, 114]. regimes [130]. Regional [102, 176, 276, 205]. regions [84, 304]. registration [290]. Regression [220, 466, 300, 278, 379, 392, 127, 29, 133, 472, 108, 252, 227, 201, 50, 279, 140, 75, 239, 296, 152, 163, 169, 238, 112, 264, 427, 471, 128, 493, 460, 342, 30, 73]. regressive [181, 467, 297, 156, 435, 395, 446, 368]. regular [262, 164]. related [288]. relation [362, 212]. relationship [128]. relationships [162]. Relative [36, 176]. reliability [9]. relief [266]. renal [387]. repeat [159, 47]. repeated [358, 38, 152, 234, 257]. replicated [449]. replication [22]. Report [126]. reported [181]. reporting [351]. request [135]. require [436]. research [8, 305, 195, 263, 250]. residential [180]. residual [279, 194]. resonance [111, 288, 203, 446, 90]. respiratory [369]. respondent [489]. respondent-driven [489]. response [182, 451, 381, 258, 139, 196, 30, 40]. responses [110, 340, 227, 296, 185]. restriction [122]. restrictive [247]. retail [417]. retesting [233]. return [334, 494]. return-inflation [494]. returns [379]. reviews [222]. ribonucleic [97, 454, 329]. richness [335]. Rio [333]. risk [236, 8, 36, 422, 143, 138, 137, 359, 349, 291, 478]. risks [211, 477, 352, 18, 459, 490, 176, 164, 16]. river [223, 169]. RNA [346]. RNASeqDesign [454]. road [141, 8]. Robust [142]. role [403]. Romance [401]. rotating [170]. rules [408]. runs [95].

Safety [214, 8, 366]. sales [272, 368]. salvage [130]. sample [48, 118, 261, 170, 20, 16]. samples [142, 492, 67]. sampling [60, 11, 425, 193, 339, 53, 489, 73]. sanitation [476]. scalar [408, 252]. scale [379, 301, 473, 106, 299]. scaled [64]. scales [217]. scenario [119]. scenario-based [119]. schedule [254]. schizophrenia [114]. school [299]. science [308]. sclerosis [261]. score [313, 178]. scores [391, 205]. scoring [192]. Scottish [143]. scrapie [82]. sea [375]. search [404]. searching [186]. second [59, 298, 328, 76]. second-harmonic [59]. second-order [328]. secondary [403]. sector [325]. Segmentation [375, 483]. seismic [251]. seizure [440]. Selection [64, 492, 423, 386, 347, 173, 441, 201, 225, 37, 42, 497, 484, 289, 266, 346, 405, 88, 131, 20, 198, 302, 456]. selective [281, 348]. Self [407, 181, 24]. self-controlled [24]. Self-exciting [407]. self-reported [181]. semantic [182]. semi [483, 495]. semi-Markov [483, 495]. semicompeting [211, 352]. semicontinuous [381]. semilocal [286]. Semiparametric [432, 168, 23, 28, 127, 187, 328, 211, 71, 128, 86, 476]. Senate [237]. sensitive [184, 40]. Sensitivity [194, 129, 324]. sensory [483]. sequence [124, 135, 187]. sequences [66]. sequencing [454, 329, 346]. sequential [214, 365, 255, 25, 114]. serial [173]. series [315, 8, 326, 52, 24, 160, 243, 321, 369]. serological [380, 199, 13]. set [72]. sets [64, 456]. settings [200]. settlement [298]. severe [235]. severity [394, 289]. sex [360]. sexually [31]. Shape [83, 7, 409, 379, 106]. shaped [7]. shared [215]. Sharing [310, 449]. shark [356]. sharp [374]. shifts [495]. short [159, 122, 235, 113, 47]. short- [235]. short-term [122]. shrinkage [77]. Signal [379, 348]. signals [377]. signature [302]. significant [72]. Similarity [286]. Similarity-based [286]. simplex [31]. simulated [337, 334]. Simulation [359, 88, 262]. simulators [266]. Simultaneous [49, 52, 394, 402, 299, 149]. Simultaneously [354, 192]. sites [463, 339]. size [358, 391, 306, 48]. skewed [388]. slab [484]. Small [290, 388]. Smooth [120, 164, 374, 323]. smoothing [82, 156, 291, 86]. snow [41]. social [481, 13, 310, 177, 309, 162, 229, 228]. socio [144]. socio-economic [144]. soft [150, 32]. software [19]. sojourn [370]. some [485, 93]. sonar [270]. sorting [111]. sounder [60]. Source [312, 28]. sources [107]. south [236, 204]. south-eastern [236]. south-west [204]. Space [96, 410, 92, 384, 145, 360, 5, 482, 229]. space-time [410]. spaceborne [127]. spaces [7]. sparse [386, 463, 116, 321].

Spatial [421, 91, 203, 363, 273, 191, 136, 419, 333, 426, 173, 156, 205, 172, 107, 405, 407, 131, 232, 10, 251]. Spatially [223, 188, 395, 392, 117, 285, 172, 446]. spatiodynamic [236]. Spatiotemporal [268, 112, 96, 334, 189, 381, 231, 291]. special [432]. species [335]. specific [450, 487, 290, 23, 377]. specification [447]. spectra [54]. spectral [316]. spectrometry [34, 155]. spectrum [301]. spike [484]. spike-and-slab [484]. spines [202]. spline [480]. splines [278, 145]. spoken [401]. spontaneous [477]. sports [134]. spraying [287]. stable [197]. stage [293, 428, 1, 196, 265, 263, 271]. stages [154]. stalked [471]. standardization [294]. Standardized [378, 205]. star [7]. star-shaped [7]. State [482, 92, 333, 145]. states [123]. stationary [419, 333, 33, 482]. Statistical [191, 330, 151, 62, 409, 496, 105, 55, 401, 348, 308]. statistics [177, 417]. status [181, 13, 477]. statuses [93]. stayer [320, 370]. stayers [325]. stem [460]. step [52, 291]. stimulating [235]. Stochastic [376, 213, 17, 404, 464, 364, 277]. stock [379, 494]. storm [276, 289]. strategic [237]. strategies [186, 1]. strategy [327]. stratification [144, 324]. Stratified [410, 178, 271]. strength [95]. stroke [252]. Structural [12, 151, 325, 87]. structure [363, 463, 229, 364]. structured [227, 55, 493, 198]. structures [172, 228]. studies [200, 278, 222, 292, 365, 23, 293, 3, 53, 221, 427, 86, 180, 246, 198, 413]. Study [311, 56, 300, 182, 358, 44, 197, 127, 148, 328, 477, 362, 454, 140, 80, 464, 261, 299, 490, 262, 58, 114, 259, 370, 458, 350, 235, 77, 247]. studying [481, 7]. subclone [455]. Subgroup [491, 302, 382, 38, 76]. subject [351, 27, 139]. Subsample [75]. suicide [195]. sum [118]. superba [60]. supervised [348]. surface [451, 243]. surgery [83]. surgical [140, 490]. surveillance [58]. survey [142, 388, 339, 184, 118, 485, 183, 170, 139]. surveys [453, 473]. Survival [362, 147, 171, 292, 56, 168, 439, 255, 328, 497, 215, 2, 232, 176, 130, 69, 460, 93, 413]. survivors [68]. Swiss [350]. switching [467]. Switzerland [58]. symptoms [399]. synchronization [377]. syndrome [44, 328, 413]. synthesizing [397]. systematic [222, 252, 107]. systemic [387]. systems [159, 52, 449, 219, 71, 312, 482, 109, 367].

table [19]. Tackling [280]. tail [142, 385]. tandem [159, 47]. target [173]. targeted [443, 436, 230]. Tariffs [313]. teenage [148]. telematics [406]. temperature [333, 122, 188]. Temporal [429, 483, 238]. tensor [261]. term [44, 122, 235, 113]. terminal [470]. termination [293]. terrorism [81]. Terry [134, 65]. test [424, 326, 118, 205, 261, 77]. Testing [121, 123, 471, 412, 101, 233, 192, 55]. tests [222, 257]. Texas [349]. texture [33]. their [148]. themed [177, 308, 431]. theoretic [436]. therapies [133, 443]. therapy [434, 230]. thermal [287]. three [151]. threshold [201, 216, 289, 153]. tidal [69]. Time [174, 380, 450, 410, 8, 136, 326, 292, 341, 52, 280, 350, 96, 384, 160, 494, 216, 120, 360, 147, 25, 490, 5, 243, 491, 321, 434, 393, 263, 413, 369]. time-dependent [350]. time-specific [450]. time-to-event [136, 490, 491]. time-to-toxicity [434]. times [182, 225, 442, 221, 147, 393, 370]. timescales [215]. timing [234]. tissue [203]. Topic [457]. total [255]. totals [453]. tournaments [134]. toxicity [255, 433, 435, 434]. traces [212]. Tracking [92]. tract [399, 108]. trade [313, 313]. traffic [141, 452]. trait [63]. traits [192, 55]. trajectories [124, 243]. transcripts [346]. transformation [458]. transformed [445]. transition [411, 179, 370]. transitions [374]. transmission [322, 174]. transmitted [31]. transplantation [460]. transportation [312]. treated [350]. Treatment [443, 292, 408, 487, 391, 465, 350, 384, 351, 216, 39, 79, 114, 1, 130, 491, 90, 161]. trees [487, 68]. tremor [321]. trends [332, 481, 429, 291]. triadic [228]. trial [255, 328, 433, 435, 294, 431, 98, 271, 76]. Trials [114, 281, 175, 382, 487, 193, 465, 115, 254, 439, 441, 48, 474, 443, 224, 27, 438, 258, 336, 76, 67, 428, 434, 196, 194, 302]. triangulation [100]. triple [438]. tropical [276]. tumour [464, 93, 45]. tumours [409]. twin [373]. twins [110]. Two [95, 428, 196, 263, 271, 36, 420, 52, 437, 293, 154, 31, 152, 261, 203, 257, 294, 1, 265]. two-agent [294]. two-component [152]. Two-level [95]. two-sample [261]. Two-stage [428, 196, 263, 271, 293, 1, 265]. two-step [52]. two-tissue [203]. type [144, 31, 234].

UK [418]. unbalanced [168]. unbiased [36]. Uncertainty [117, 488, 9, 338, 12, 359, 289, 87]. uncertainty-directed [488]. undergoing [79]. underwater [270]. unemployment [459]. unexploded [137]. unified [433]. uniform [489]. unit [10]. units [309]. univariate [378]. unknown [408, 132, 150, 93]. unmeasured [292, 468]. unobserved [325, 179]. Unravelling [406]. unrepresentative [178]. unrest [268]. uptake [45]. Urban [349]. urinary [399]. use [411, 253, 337, 183, 170]. Using [19, 385, 58, 322, 15, 423, 101, 100, 396, 34, 26, 453, 214, 60, 316, 496, 298, 61, 22, 199, 306, 252, 226, 418, 255, 190, 441, 102, 217, 486, 3, 156, 404, 329, 137, 215, 324, 317, 484, 349, 435, 210, 452, 480, 485, 176, 229, 67, 128, 164, 65, 428, 47, 170, 245, 1, 434, 412]. utility [439]. utility-based [439]. utilizing Vaccine [214]. Validation [39, 293]. validatory [289]. value [289]. valued [326, 467]. values [398, 78, 213]. variability [495, 190]. Variable [200, 201, 88, 386, 411, 148, 441, 497, 468, 484, 185, 405, 131, 128, 198, 456]. variable-selection [456]. variables [399, 391, 388, 64, 226, 144, 96, 445, 493, 139]. variance [322, 348, 470]. variant [46]. variation [269, 176]. variety [67]. varying [380, 280, 494, 258, 172, 147, 395, 446, 174, 263, 413]. varying-coefficient [258]. vector [368]. vehicle [394]. vehicles [366]. velocity [41]. ventilation [69]. ventricle [496]. verbal [79]. versatile [103]. versus [57]. Very [113]. Very-short-term [113]. veterinary [58]. via [449, 111, 445, 493, 90, 265, 161]. videos [270]. vine [262]. viral [22]. virtualized [71]. virus [322, 44, 463, 253, 97, 216, 31, 395, 413]. Visualizing [9]. vital [290]. volatility [315, 374, 160]. volcanic [321]. volume [85, 125, 166, 165, 209, 241, 240, 283, 282, 344, 343, 415, 414, 499, 498, 69]. voting [116].

wage [267]. War [137]. wavelet [33, 256]. wavelet-based [256]. wavelets [217]. weak [452]. weather [21, 131]. Weibull [219, 448]. weighing [274]. weight [47]. Weighted [37, 398, 223]. Weighting [53]. weights [170]. wellbeing [450]. west [204]. Western [204]. wet [5]. whale [356]. while [107]. who [489]. whole [316]. whole-brain [316]. Wilcoxon [118]. win [272]. wind [113]. windows [393]. Within [24, 158]. within-household [158]. Within-individual [24]. without [365]. women [79, 128]. working [427]. Workload [71]. World [137].

xenography [464]. years [91]. yeast [260].

Zelterman [356]. zero [365, 354, 239, 263]. zero-inflated [365, 239, 263]. zones [137].

References

Wahed:2010:ITS

[1] Abdus S. Wahed. Inference for two-stage adaptive treatment strategies using mixture distributions. Applied Statistics, 59 (1):1–18, January 2010. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Morita:2010:BHM

[2] Satoshi Morita, Peter F. Thall, B. Nebiyou Bekele, and Paul Mathew. A Bayesian hierarchical mixture model for platelet-derived growth factor receptor phosphorylation to improve estimation of progression-free survival in prostate cancer. Applied Statistics, 59 (1):19–34, January 2010. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lambert:2010:ECF

[3] P. C. Lambert, P. W. Dickman, C. L. Weston, and J. R. Thompson. Estimating the cure fraction in population-based cancer studies by using finite mixture models. Applied Statistics, 59(1):35– 55, January 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

OConnor:2010:FPC

[4] Emma O’Connor, Nick Fieller, Andrew Holmes, John C. Waterton, and Edward Ainscow. Functional principal component analyses of biomedical images as outcome measures. Applied Statistics, 59 (1):57–76, January 2010. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sahu:2010:FPA

[5] Sujit K. Sahu, Alan E. Gelfand, and David M. Holland. Fusing point and areal level space–time data with application to wet deposition. Applied Statistics, 59(1):77–103, January 2010. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wei:2010:NBG

[6] Peng Wei and Wei Pan. Networkbased genomic discovery: application and comparison of Markov random-field models. Applied Statistics, 59(1):105– 125, January 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hotz:2010:SSP

[7] T. Hotz, S. Huckemann, A. Munk, D. Gaffrey, and B. Sloboda. Shape spaces for prealigned star-shaped objects—studying the growth of plants by principal components analysis. Applied Statistics, 59(1):127–143, January 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bijleveld:2010:MNL

[8] Frits Bijleveld, Jacques Commandeur, Siem Jan Koopman, and Kees van Montfort. Multivariate non-linear time series modelling of exposure and risk in road safety research. Applied Statistics, 59(1): 145–161, January 2010. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ebden:2010:VUR

[9] Mark Ebden, Armin Stranjak, and Stephen Roberts. Visualizing uncertainty in reliability functions with application to aero engine overhaul. Applied Statistics, 59(1):163–173, January 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tassone:2010:DSM

[10] Eric C. Tassone, Marie Lynn Miranda, and Alan E. Gelfand. Disaggregated spatial modelling for areal unit categorical data. Applied Statistics, 59(1):175– 190, January 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Diggle:2010:GIU

[11] Peter J. Diggle, Raquel Menezes, and Ting li Su. Geostatistical inference under preferential sampling. Applied Statistics, 59(2):191–232, March 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jackson:2010:SPU

[12] Christopher H. Jackson, Linda D. Sharples, and Simon G. Thompson. Structural and parameter uncertainty in Bayesian cost-effectiveness models. Applied Statistics, 59(2):233–253, March 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Goeyvaerts:2010:EID

[13] Nele Goeyvaerts, Niel Hens, Benson Ogunjimi, Marc Aerts, Ziv Shkedy, Pierre Van Damme, and Philippe Beutels. Estimating infectious disease parameters from data on social contacts and serological status. Applied Statistics, 59(2):255–277, March 2010. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Leith:2010:FIC

[14] Nadja A. Leith and Richard E. Chandler. A framework for interpreting climate model outputs. Applied Statistics, 59(2):279–296, March 2010. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Aston:2010:LPA

[15] John A. D. Aston, Jeng-Min Chiou, and Jonathan P. Evans. Linguistic pitch analysis using functional principal component mixed effect models. Applied Statistics, 59(2):297–317, March 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zhou:2010:RDR

[16] Yijie Zhou, Francesca Dominici, and Thomas A. Louis. Racial disparities in risks of mortality in a sample of the US Medicare population. Applied Statistics, 59(2):319–339, March 2010. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gillespie:2010:BIG

[17] Colin S. Gillespie and Andrew Golightly. Bayesian inference for generalized stochastic population growth models with application to aphids. Applied Statistics, 59(2):341–357, March 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lo:2010:CMD

[18] Simon M. S. Lo and Ralf A. Wilke. A copula model for dependent competing risks. Applied Statistics, 59(2):359– 376, March 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Altham:2010:URD

[19] P. M. E. Altham and Robin K. S. Hankin. Using recently developed software on a 2 × 2 table of matched pairs with incompletely classified data. Applied Statistics, 59(2):377–379, March 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wiesenfarth:2010:BGS

[20] Manuel Wiesenfarth and Thomas Kneib. Bayesian geoadditive sample selection models. Applied Statistics, 59(3):381– 404, May 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

DiNarzo:2010:BHA

[21] A. F. Di Narzo and D. Cocchi. A Bayesian hierarchical approach to ensemble weather forecasting. Applied Statistics, 59(3):405–422, May 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Follmann:2010:EVI

[22] Dean Follmann, Jing Qin, and Yo Hoshino. Estimation of viral infection and replication in cells by using convolution models. Applied Statistics, 59(3):423–435, May 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Huang:2010:SME

[23] Y. Huang and M. S. Pepe. Semiparametric methods for evaluating the covariate-specific predictiveness of continuous markers in matched case–control studies. Applied Statistics, 59(3):437– 456, May 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Farrington:2010:WID

[24] C. Paddy Farrington and Mounia N. Hocine. Within-individual dependence in self-controlled case series models for recurrent events. Applied Statistics, 59 (3):457–475, May 2010. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Robinson:2010:CTB

[25] G. K. Robinson. Continuous time Brownian motion models for analysis of sequential data. Applied Statistics, 59(3): 477–494, May 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bruckers:2010:LCA

[26] Liesbeth Bruckers, Jan Serroyen, Geert Molenberghs, Herman Slaets, and Willem Goeyvaerts. Latent class analysis of persistent disturbing behaviour patients by using longitudinal profiles. Applied Statistics, 59(3):495–512, May 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Long:2010:ECE

[27] Qi Long, Roderick J. A. Little, and Xihong Lin. Estimating causal effects in trials involving multitreatment arms subject to non-compliance: a Bayesian framework. Applied Statistics, 59(3): 513–531, May 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rodrigues:2010:SAP

[28] Alexandre Rodrigues, Peter Diggle, and Renato Assuncao. Semiparametric approach to point source modelling in epidemiology and criminology. Applied Statistics, 59(3):533–542, May 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gasparini:2010:BDF

[29] Mauro Gasparini, Stuart Bailey, and Beat Neuenschwander. Bayesian dose finding in oncology for drug combinations by copula regression. Applied Statistics, 59(3):543–544, May 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Yin:2010:ARB

[30] Guosheng Yin and Ying Yuan. Authors’ response: Bayesian dose finding in oncology for drug combinations by copula regression. Applied Statistics, 59(3): 544–546, May 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mahiane:2010:MMC

[31] S. Guy Mahiane, Eugène-P. Ndong Nguéma, Carel Pretorius, and Bertran Auvert. Mathematical models for coinfection by two sexually transmitted agents: the human immunodeficiency virus and herpes simplex virus type 2 case. Applied Statistics, 59(4):547– 572, August 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rudoy:2010:BCP

[32] Daniel Rudoy, Shelten G. Yuen, Robert D. Howe, and Patrick J.

Wolfe. Bayesian change-point analysis for atomic force microscopy and soft material indentation. Applied Statistics, 59 (4):573–593, August 2010. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Eckley:2010:LSW

[33] Idris A. Eckley, Guy P. Nason, and Robert L. Treloar. Locally stationary wavelet fields with application to the modelling and analysis of image texture. Applied Statistics, 59(4):595– 616, August 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Browne:2010:MEM

[34] William J. Browne, Ian L. Dryden, Kelly Handley, Shahid Mian, and Dirk Schadendorf. Mixed effect modelling of proteomic mass spectrometry data by using Gaussian mixtures. Applied Statistics, 59(4):617–633, August 2010. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zhang:2010:BMB

[35] Jian Zhang. A Bayesian model for biclustering with applications. Applied Statistics, 59(4):635–656, August 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Carter:2010:RRE

[36] Rickey E. Carter, Yan Lin, Stuart R. Lipsitz, Robert G. Newcombe, and Kathie L. Hermayer. Relative risk estimated from the ratio of two median unbiased estimates. Applied Statistics, 59 (4):657–671, August 2010. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Li:2010:WAU

[37] Jialiang Li and Jason P. Fine. Weighted area under the receiver operating characteristic curve and its application to gene selection. Applied Statistics, 59(4): 673–692, August 2010. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Li:2010:ICD

[38] Jialiang Li and Shuangge Ma. Intervalcensored data with repeated measurements and a cured subgroup. Applied Statistics, 59(4):693–705, August 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nielsen:2010:VMI

[39] Lars Hougaard Nielsen and Niels Keiding. Validation of methods for identifying discontinuation of treatment from prescription data. Applied Statistics, 59 (4):707–722, August 2010. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

vandenHout:2010:EPS

[40] Ardo van den Hout, Ulf Böckenholt, and Peter G. M. Van Der Heijden. Estimating the prevalence of sensitive behaviour and cheating with a dual design for direct questioning and randomized response. Applied Statistics, 59(4):723– 736, August 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rougier:2010:PSV

[41] Jonathan Rougier and Martin Kern. Predicting snow velocity in large chute flows under different environmental conditions. Applied Statistics, 59(5):737– 760, November 2010. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Loh:2010:ECM

[42] Ji Meng Loh and Woncheol Jang. Estimating a cosmological mass bias parameter with bootstrap bandwidth selection. Applied Statistics, 59(5):761– 779, November 2010. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jank:2010:FME

[43] Wolfgang Jank, Galit Shmueli, and Shu Zhang. A flexible model for estimating price dynamics in on-line auctions. Applied Statistics, 59(5):781–804, November 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chen:2010:MLT

[44] Jianwei Chen. Modelling long-term human immunodeficiency virus dynamic models with application to acquired immune deficiency syndrome clinical study. Applied Statistics, 59(5):805–820, November 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zhang:2010:LIA

[45] Xiaoxi Zhang, Timothy D. Johnson, Roderick J. A. Little, and Yue Cao. Longitudinal image analysis of tumour– healthy brain change in contrast uptake induced by radiation. Applied Statistics, 59(5):821–838, November 2010. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chadeau-Hyam:2010:AHM

[46] Marc Chadeau-Hyam, Paul S. Clarke, Chantal Guihenneuc-Jouyaux, Simon N. Cousens, Robert G. Will, and Azra C. Ghani. An application of hidden Markov models to the French variant Creutzfeldt–Jakob disease epidemic. Applied Statistics, 59(5):839–853, November 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tvedebrink:2010:EWE

[47] Torben Tvedebrink, Poul Svante Eriksen, Helle Smidt Mogensen, and Niels Morling. Evaluating the weight of evidence by using quantitative short tandem repeat data in DNA mixtures. Applied Statistics, 59(5):855–874, November 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kikuchi:2010:BBA

[48] Takashi Kikuchi and John Gittins. A behavioural Bayes approach for sample size determination in cluster randomized clinical trials. Applied Statistics, 59(5): 875–888, November 2010. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Azais:2010:SCB

[49] J. M. Azaïs, S. Bercu, J. C. Fort, A. Lagnoux, and P. Lé. Simultaneous confidence bands in curve prediction applied to load curves. Applied Statistics, 59(5):889–904, November 2010. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lee:2010:BQR

[50] Duncan Lee and Tereza Neocleous. Bayesian quantile regression for count data with application to environmental epidemiology. Applied Statistics, 59(5): 905–920, November 2010. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2010:IA

[51] Anonymous. Index of authors. Applied Statistics, 59(5):921–924, November 2010. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

DiFonzo:2011:STS

[52] Tommaso Di Fonzo and Marco Marini. Simultaneous and two-step reconciliation of systems of time series: methodological and practical issues. Applied Statistics, 60(2):143–164, March 2011. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Li:2011:WMP

[53] Yan Li, Barry I. Graubard, and Ralph DiGaetano. Weighting methods for population-based case–control studies with complex sampling. Applied Statistics, 60(2):165–185, March 2011. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zhong:2011:BMD

[54] Mingjun Zhong, Mark Girolami, Karen Faulds, and Duncan Graham. Bayesian methods to detect dye-labelled DNA oligonucleotides in multiplexed Raman spectra. Applied Statistics, 60(2):187– 206, March 2011. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Li:2011:BNP

[55] Meijuan Li and Timothy Hanson. Bayesian non-parametric multivariate statistical models for testing association between quantitative traits and candidate genes in structured populations. Applied Statistics, 60(2):207–219, March 2011. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Graham:2011:JMS

[56] Petra L. Graham, Louise M. Ryan, and Mary A. Luszcz. Joint modelling of survival and cognitive decline in the Australian Longitudinal Study of Ageing. Applied Statistics, 60(2):221–238, March 2011. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

deBlasio:2011:FEN

[57] Birgitte Freiesleben de Blasio, Taral Guldahl Seierstad, and Odd O. Aalen. Frailty effects in networks: comparison and identification of individual heterogeneity versus preferential attachment in evolving networks. Applied Statistics, 60 (2):239–259, March 2011. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Schrodle:2011:UIN

[58] Birgit Schrödle, Leonhard Held, Andrea Riebler, and Jürg Danuser. Using integrated nested Laplace approximations for the evaluation of veterinary surveillance data from Switzerland: a casestudy. Applied Statistics, 60(2):261– 279, March 2011. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Biedermann:2011:ODG

[59] Stefanie Biedermann and David C. Woods. Optimal designs for generalized non-linear models with application to second-harmonic generation experiments. Applied Statistics, 60(2):281– 299, March 2011. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cox:2011:EDA

[60] Martin J. Cox, David L. Borchers, David A. Demer, George R. Cutter, and Andrew S. Brierley. Estimating the density of Antarctic krill (Euphausia superba) from multi-beam echo-sounder observations using distance sampling methods. Applied Statistics, 60(2):301– 316, March 2011. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Drovandi:2011:ABC

[61] Christopher C. Drovandi, Anthony N. Pettitt, and Malcolm J. Faddy. Approximate Bayesian computation using indirect inference. Applied Statistics, 60(3): 317–337, May 2011. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Maul:2011:SMG

[62] Philip R. Maul, Peter C. Robinson, and Paul J. Northrop. Statistical modelling of graphite brick cracking in advanced gas-cooled reactors. Applied Statistics, 60(3):339–353, May 2011. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Au:2011:MME

[63] Kinman Au, Rongheng Lin, and Andrea S. Foulkes. Mixture modelling as an exploratory framework for genotype– trait associations. Applied Statistics, 60 (3):355–375, May 2011. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gertheiss:2011:SOS

[64] Jan Gertheiss, Sara Hogger, Cornelia Oberhauser, and Gerhard Tutz. Selection of ordinally scaled independent variables with applications to international classification of functioning core sets. Applied Statistics, 60(3):377–395, May 2011. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stern:2011:MPC

[65] Steven E. Stern. Moderated paired comparisons: a generalized Bradley– Terry model for continuous data using a discontinuous penalized likelihood function. Applied Statistics, 60(3):397– 415, May 2011. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Scott:2011:AMC

[66] Marc A. Scott. Affinity models for career sequences. Applied Statistics, 60(3): 417–436, May 2011. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smith:2011:DAV

[67] Alison B. Smith, Robin Thompson, David G. Butler, and Brian R. Cullis. The design and analysis of variety trials using mixtures of composite and individual plot samples. Applied Statistics, 60 (3):437–455, May 2011. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Su:2011:ITE

[68] Xiaogang Su, Karen Meneses, Patrick McNees, and Wesley O. Johnson. Interaction trees: exploring the differential effects of an intervention programme for breast cancer survivors. Applied Statistics, 60(3):457–474, May 2011. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wang:2011:ECE

[69] Weiwei Wang, Daniel Scharfstein, Chenguang Wang, Michael Daniels, Dale Needham, and Roy Brower. Estimating the causal effect of low tidal volume ventilation on survival in patients with acute lung injury. Applied Statistics, 60 (4):475–496, August 2011. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bisgaard:2011:QEP

[70] Søren Bisgaard and Davit Khachatryan. Quasi-experiments on process dynamics. Applied Statistics, 60(4):497– 517, August 2011. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Li:2011:WPO

[71] Ta-Hsin Li. Workload portfolio optimization for virtualized computer systems based on semiparametric quantile function estimation. Applied Statistics, 60(4):519–539, August 2011. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shahbaba:2011:BGS

[72] Babak Shahbaba, Robert Tibshirani, Catherine M. Shachaf, and Sylvia K. Plevritis. Bayesian gene set analysis for identifying significant biological pathways. Applied Statistics, 60(4):541– 557, August 2011. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zhou:2011:PLR

[73] Haibo Zhou, Jinhong You, Guoyou Qin, and Matthew P. Longnecker. A partially linear regression model for data from an outcome-dependent sampling design. Applied Statistics, 60(4):559– 574, August 2011. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Paddock:2011:PBE

[74] Susan M. Paddock and Thomas A. Louis. Percentile-based empirical distribution function estimates for performance evaluation of healthcare providers. Applied Statistics, 60(4):575– 589, August 2011. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Little:2011:SIL

[75] Roderick J. Little and Nanhua Zhang. Subsample ignorable likelihood for regression analysis with missing data. Applied Statistics, 60(4):591–605, August 2011. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Scharfstein:2011:ASE

[76] Daniel Scharfstein, Georgiana Onicescu, Steven Goodman, and Rachel Whitaker. Analysis of subgroup effects in randomized trials when subgroup membership is missing: application to the second Multicenter Automatic Defibrillator Intervention Trial. Applied Statistics, 60 (4):607–617, August 2011. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Satagopan:2011:PPT

[77] Jaya M. Satagopan, Qin Zhou, Susan A. Oliveria, Stephen W. Dusza, Martin A. Weinstock, Marianne Berwick, and Allan C. Halpern. Properties of preliminary test estimators and shrinkage estimators for evaluating multiple exposures — application to questionnaire data from the ‘Study of nevi in children’. Applied Statistics, 60(4):619– 632, August 2011. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Huang:2011:BIA

[78] Ying Huang, Youyi Fong, John Wei, and Ziding Feng. Borrowing information across populations in estimating positive and negative predictive values. Applied Statistics, 60(5):633–653, November 2011. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rolfe:2011:BEE

[79] Margaret I. Rolfe, Kerrie L. Mengersen, Katharine J. Vearncombe, Brooke Andrew, and Geoffrey F. Beadle. Bayesian estimation of extent of recovery for aspects of verbal memory in women undergoing adjuvant chemotherapy treatment for breast cancer. Applied Statistics, 60(5):655–674, November 2011. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

OKeeffe:2011:CSC

[80] Aidan G. O’Keeffe, Brian D. M. Tom, and Vernon T. Farewell. A case-study in the clinical epidemiology of psoriatic arthritis: multistate models and causal arguments. Applied Statistics, 60(5): 675–699, November 2011. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kyung:2011:NFT

[81] Minjung Kyung, Jeff Gill, and George Casella. New findings from terrorism data: Dirichlet process random-effects models for latent groups. Applied Statistics, 60(5):701–721, November 2011. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bohning:2011:CRE

[82] Dankmar Böhning, Ronny Kuhnert, and Victor Del Rio Vilas. Capture– recapture estimation by means of empirical Bayesian smoothing with an application to the geographical distribution of hidden scrapie in Great Britain. Applied Statistics, 60(5):723– 741, November 2011. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Faraway:2011:SCA

[83] Julian J. Faraway and Carroll-Ann Trotman. Shape change along geodesics with application to cleft lip surgery. Applied Statistics, 60(5):743–755, November 2011. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chakraborty:2011:PPM

[84] Avishek Chakraborty, Alan E. Gelfand, Adam M. Wilson, Andrew M. Latimer, and John A. Silander. Point pattern modelling for degraded presence-only data over large regions. Applied Statistics, 60(5):757–776, November 2011.

CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2011:IAV

[85] Anonymous. Index of authors, volume 60, 2011. Applied Statistics, 60(5):777– 780, November 2011. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wang:2012:FSA

[86] Yuanjia Wang, Chiahui Huang, Yixin Fang, Qiong Yang, and Runze Li. Flexible semiparametric analysis of longitudinal genetic studies by reduced rank smoothing. Applied Statistics, 61(1):1– 24, January 2012. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Strong:2012:MSU

[87] Mark Strong, Jeremy E. Oakley, and Jim Chilcott. Managing structural uncertainty in health economic decision models: a discrepancy approach. Applied Statistics, 61(1):25–45, January 2012. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Reich:2012:VSH

[88] Brian J. Reich, Eric Kalendra, Curtis B. Storlie, Howard D. Bondell, and Montserrat Fuentes. Variable selection for high dimensional Bayesian density estimation: application to human exposure simulation. Applied Statistics, 61 (1):47–66, January 2012. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Yu:2012:AAM

[89] Binbing Yu and Chuan Zhou. Assessing the accuracy of a multiphase diagnosis procedure for dementia. Applied Statistics, 61(1):67–81, January 2012. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wu:2012:PTE

[90] Jincao Wu, Timothy D. Johnson, Craig J. Galbán, Thomas L. Chenevert, Charles R. Meyer, Alnawaz Rehemtulla, Daniel A. Hamstra, and Brian D. Ross. Predicting treatment efficacy via quantitative magnetic resonance imaging: a Bayesian joint model. Applied Statistics, 61(1):83–98, January 2012. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Li:2012:SML

[91] Ye Li, Patrick Brown, Håvard Rue, Mustafa al Maini, and Paul Fortin. Spatial modelling of lupus incidence over 40 years with changes in census areas. Applied Statistics, 61(1):99–115, January 2012. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chen:2012:TMI

[92] Shi Chen, John Fricks, and Matthew J. Ferrari. Tracking measles infection through non-linear state space models. Applied Statistics, 61(1):117–134, January 2012. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Yuan:2012:EPF

[93] Ying Yuan, Peter F. Thall, and Johannes E. Wolff. Estimating progression-free survival in paediatric brain tumour patients when some progression statuses are unknown. Applied Statistics, 61(1):135–149, January 2012.

CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Besbeas:2012:KFI

[94] Panagiotis Besbeas and Byron J. T. Morgan. Kalman filter initialization for integrated population modelling. Applied Statistics, 61(1):151–162, January 2012. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Schoen:2012:TLD

[95] Eric D. Schoen and Robert W. Mee. Two-level designs of strength 3 and up to 48 runs. Applied Statistics, 61(1):163– 174, January 2012. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ippoliti:2012:STM

[96] L. Ippoliti, P. Valentini, and D. Gamerman. Space–time modelling of coupled spatiotemporal environmental variables. Applied Statistics, 61(2):175–200, March 2012. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Grun:2012:MHI

[97] Bettina Grün and Kurt Hornik. Modelling human immunodeficiency virus ribonucleic acid levels with finite mixtures for censored longitudinal data. Applied Statistics, 61(2):201–218, March 2012. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Yin:2012:PIT

[98] Guosheng Yin, Nan Chen, and J. Jack Lee. Phase II trial design with Bayesian adaptive randomization and predictive probability. Applied Statistics, 61(2): 219–235, March 2012. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gilmour:2012:BOE

[99] Steven G. Gilmour and Luzia A. Trinca. Bayesian L-optimal exact design of experiments for biological kinetic models. Applied Statistics, 61(2):237–251, March 2012. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bray:2012:QND

[100] David J. Bray, Steven G. Gilmour, Felicity J. Guild, and Ambrose C. Taylor. Quantifying nanoparticle dispersion by using the area disorder of Delaunay triangulation. Applied Statistics, 61 (2):253–275, March 2012. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Black:2012:GTH

[101] Michael S. Black, Christopher R. Bilder, and Joshua M. Tebbs. Group testing in heterogeneous populations by using halving algorithms. Applied Statistics, 61(2):277–290, March 2012. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kang:2012:CON

[102] Emily L. Kang, Noel Cressie, and Stephan R. Sain. Combining outputs from the North American Regional Climate Change Assessment Program by using a Bayesian hierarchical model. Applied Statistics, 61(2):291–313, March 2012. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pipper:2012:VMC

[103] Christian Bressen Pipper, Christian Ritz, and Hans Bisgaard. A versatile method for confirmatory evaluation of the effects of a covariate in multiple models. Applied Statistics, 61(2):315– 326, March 2012. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Andersen:2012:AFE

[104] Stina W. Andersen and George C. Runger. Automated feature extraction from profiles with application to a batch fermentation process. Applied Statistics, 61(2):327–344, March 2012. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gilmour:2012:ODE

[105] Steven G. Gilmour and Luzia A. Trinca. Optimum design of experiments for statistical inference. Applied Statistics, 61 (3):345–401, May 2012. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mayr:2012:GAM

[106] Andreas Mayr, Nora Fenske, Benjamin Hofner, Thomas Kneib, and Matthias Schmid. Generalized additive models for location, scale and shape for high dimensional data—a flexible approach based on boosting. Applied Statistics, 61(3): 403–427, May 2012. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Paciorek:2012:CSI

[107] Christopher J. Paciorek. Combining spatial information sources while accounting for systematic errors in proxies. Applied Statistics, 61(3):429–451, May 2012. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Goldsmith:2012:LPF

[108] Jeff Goldsmith, Ciprian M. Crainiceanu, Brian Caffo, and Daniel Reich. Longitudinal penalized functional regression for cognitive outcomes on neuronal tract measurements. Applied Statistics, 61(3): 453–469, May 2012. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sun:2012:EPR

[109] Mingxuan Sun, Guy Lebanon, and Paul Kidwell. Estimating probabilities in recommendation systems. Applied Statistics, 61(3):471–492, May 2012. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ekholm:2012:LCM

[110] Anders Ekholm, Jukka Jokinen, John W.

McDonald, and Peter W. F. Smith. A latent class model for bivariate binary responses from twins. Applied Statistics, 61(3):493–514, May 2012. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kao:2012:CMD

[111] Ming-Hung Kao, Abhyuday Mandal, and John Stufken. Constrained multiobjective designs for functional magnetic resonance imaging experiments via a modified non-dominated sorting genetic algorithm. Applied Statistics, 61 (4):515–534, August 2012. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Reich:2012:SQR

[112] Brian J. Reich. Spatiotemporal quantile regression for detecting distributional changes in environmental processes. Applied Statistics, 61(4):535–553, August 2012. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pinson:2012:VST

[113] P. Pinson. Very-short-term probabilistic forecasting of wind power with generalized logit-normal distributions. Applied Statistics, 61(4):555–576, August 2012. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shortreed:2012:EOD

[114] Susan M. Shortreed and Erica E. M. Moodie. Estimating the optimal dynamic antipsychotic treatment regime: evidence from the sequential multipleassignment randomized Clinical Antipsychotic Trials of Intervention and Effectiveness schizophrenia study. Applied Statistics, 61(4):577–599, August 2012. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Friede:2012:BCM

[115] Tim Friede and Frank Miller. Blinded continuous monitoring of nuisance parameters in clinical trials. Applied Statistics, 61(4):601–618, August 2012. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hahn:2012:SFA

[116] P. Richard Hahn, Carlos M. Carvalho, and James G. Scott. A sparse factor analytic probit model for Congressional voting patterns. Applied Statistics, 61 (4):619–635, August 2012. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Foster:2012:USP

[117] Scott D. Foster, Hideyasu Shimadzu, and Ross Darnell. Uncertainty in spatially predicted covariates: is it ignorable? Applied Statistics, 61(4):637– 652, August 2012. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Natarajan:2012:EWR

[118] Sundar Natarajan, Stuart R. Lipsitz, Garrett M. Fitzmaurice, Debajyoti Sinha, Joseph G. Ibrahim, Jennifer Haas, and Walid Gellad. An extension of the Wilcoxon rank sum test for complex sample survey data. Applied Statistics, 61(4):653–664, August 2012. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Williamson:2012:FLA

[119] Daniel Williamson, Michael Goldstein, and Adam Blaker. Fast linked analyses for scenario-based hierarchies. Applied Statistics, 61(5):665–691, November 2012. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Miller:2012:SPC

[120] Claire Miller and Adrian Bowman.

Smooth principal components for investigating changes in covariances over time. Applied Statistics, 61(5):693–714, November 2012. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gromenko:2012:TEM

[121] Oleksandr Gromenko and Piotr Kokoszka.

Testing the equality of mean functions of ionospheric critical frequency curves. Applied Statistics, 61(5):715– 731, November 2012. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Golightly:2012:ELO

[122] Andrew Golightly, Richard J. Boys, Kerry M. Cameron, and Thomas von Zglinicki. The effect of late onset, short-term caloric restriction on the core temperature and physical activity in mice. Applied Statistics, 61(5):733–751, November 2012. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jupp:2012:TQS

[123] Peter E. Jupp, Peter T. Kim, Ja-Yong Koo, and Aron Pasieka. Testing quantum states for purity. Applied Statistics, 61(5):753–763, November 2012. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Barban:2012:CLC

[124] Nicola Barban and Francesco C. Billari. Classifying life course trajectories: a comparison of latent class and sequence analysis. Applied Statistics, 61(5):765–784, November 2012. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2012:IAV

[125] Anonymous. Index of authors, volume 61, 2012. Applied Statistics, 61(5):785– 787, November 2012. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chandler:2013:RER

[126] R. E. Chandler and M. S. Ridout. Report of the Editors. Applied Statistics, 62(1):1–2, January 2013. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Farah:2013:ASB

[127] Marian Farah, Athanasios Kottas, and Robin D. Morris. An application of semiparametric Bayesian isotonic regression to the study of radiation effects in spaceborne microelectronics. Applied Statistics, 62(1):3–24, January 2013. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sobotka:2013:ERB

[128] Fabian Sobotka, Rosalba Radice, Giampiero Marra, and Thomas Kneib. Estimating the relationship between women’s education and fertility in Botswana by using an instrumental variable approach to semiparametric expectile regression. Applied Statistics, 62 (1):25–45, January 2013. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Copas:2013:LBS

[129] John B. Copas. A likelihood-based sensitivity analysis for publication bias in meta-analysis. Applied Statistics, 62 (1):47–66, January 2013. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wahed:2013:EJE

[130] Abdus S. Wahed and Peter F. Thall.

Evaluating joint effects of induction– salvage treatment regimes on overall survival in acute leukaemia. Applied Statistics, 62(1):67–83, January 2013. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Scheel:2013:BHM

[131] Ida Scheel, Egil Ferkingstad, Arnoldo Frigessi, Ola Haug, Mikkel Hinnerichsen, and Elisabeth Meze-Hausken. A Bayesian hierarchical model with spatial variable selection: the effect of weather on insurance claims. Applied Statistics, 62(1):85–100, January 2013. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

McCrea:2013:ADM

[132] Rachel S. McCrea, Byron J. T. Morgan, and Diana J. Cole. Age-dependent mixture models for recovery data on animals marked at unknown age. Applied Statistics, 62(1):101–113, January 2013. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gasparini:2013:GCM

[133] Mauro Gasparini. General classes of multiple binary regression models in dose finding problems for combination therapies. Applied Statistics, 62(1):115– 133, January 2013. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cattelan:2013:DBT

[134] Manuela Cattelan, Cristiano Varin, and David Firth. Dynamic Bradley–Terry modelling of sports tournaments. Applied Statistics, 62(1):135–150, January 2013. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Durand:2013:OPC

[135] Jean-Baptiste Durand, Stéphane Girard, Victor Ciriza, and Laurent Donini. Optimization of power consumption and device availability based on point process modelling of the request sequence. Applied Statistics, 62(2):151–165, March 2013. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chang:2013:STE

[136] Howard H. Chang, Brian J. Reich, and Marie Lynn Miranda. A spatial timeto-event approach for estimating associations between air pollution and preterm birth. Applied Statistics, 62(2):167– 179, March 2013. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mahling:2013:DHR

[137] Monia Mahling, Michael Höhle, and Helmut Küchenhoff. Determining highrisk zones for unexploded World War II bombs by using point process methodology. Applied Statistics, 62(2):181– 199, March 2013. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kolamunnage-Dona:2013:MMA

[138] Ruwanthi Kolamunnage-Dona, L. Vitone, W. Greenhalf, R. Henderson, and P. R. Williamson. A multistate modelling approach for pancreatic cancer development in genetically high-risk families. Applied Statistics, 62(2):201– 212, March 2013. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

West:2013:NRA

[139] Brady T. West and Roderick J. A. Little. Non-response adjustment of survey estimates based on auxiliary variables subject to error. Applied Statistics, 62 (2):213–231, March 2013. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lipsitz:2013:BCP

[140] Stuart R. Lipsitz, Garrett M. Fitzmaurice, Scott E. Regenbogen, Debajyoti Sinha, Joseph G. Ibrahim, and Atul A. Gawande. Bias correction for the proportional odds logistic regression model with application to a study of surgical complications. Applied Statistics, 62 (2):233–250, March 2013. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anacleto:2013:MFR

[141] Osvaldo Anacleto, Catriona Queen, and Casper J. Albers. Multivariate forecasting of road traffic flows in the presence of heteroscedasticity and measurement errors. Applied Statistics, 62(2):251– 270, March 2013. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Alfons:2013:REE

[142] Andreas Alfons, Matthias Templ, and Peter Filzmoser. Robust estimation of economic indicators from survey samples based on Pareto tail modelling. Applied Statistics, 62(2):271–286, March 2013. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Finazzi:2013:MBF

[143] Francesco Finazzi, E. Marian Scott, and Alessandro Fassò. A model-based framework for air quality indices and population risk evaluation, with an application to the analysis of Scottish air quality data. Applied Statistics, 62(2):287– 308, March 2013. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic).

Hennig:2013:HFA

[144] Christian Hennig and Tim F. Liao.

How to find an appropriate clustering for mixed-type variables with application to socio-economic stratification. Applied Statistics, 62(3):309–369, May 2013. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Quenneville:2013:CIS

[145] B. Quenneville, F. Picard, and S. Fortier. Calendarization with interpolating splines and state space models. Applied Statistics, 62(3):371–399, May 2013. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ieva:2013:MFC

[146] Francesca Ieva, Anna M. Paganoni, Davide Pigoli, and Valeria Vitelli. Multivariate functional clustering for the morphological analysis of electrocardiograph curves. Applied Statistics, 62(3):401– 418, May 2013. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rathbun:2013:SAT

[147] Stephen L. Rathbun, Xiao Song, Benjamin Neustifter, and Saul Shiffman. Survival analysis with time varying covariates measured at random times by design. Applied Statistics, 62(3):419– 434, May 2013. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jackson:2013:OLV

[148] John C. Jackson, Paul S. Albert, Zhiwei Zhang, and Bruce Simons-Morton. Ordinal latent variable models and their application in the study of newly licensed teenage drivers. Applied Statistics, 62 (3):435–450, May 2013. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bobb:2013:RHM

[149] Jennifer F. Bobb, Francesca Dominici, and Roger D. Peng. Reduced hierarchical models with application to estimating health effects of simultaneous exposure to multiple pollutants. Applied Statistics, 62(3):451–472, May 2013. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Michalak:2013:AIC

[150] Sarah E. Michalak, Michael S. Hamada, and Nicolas W. Hengartner. Analysis of interval-censored data with random unknown end points: an application to soft error rate estimation. Applied Statistics, 62(3):473–486, May 2013. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mardia:2013:SAT

[151] Kanti V. Mardia. Statistical approaches to three key challenges in protein structural bioinformatics. Applied Statistics, 62(3):487–514, May 2013. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

McMillan:2013:TCC

[152] Garnett P. McMillan, Timothy E. Hanson, Gabrielle Saunders, and Frederick J. Gallun. A two-component circular regression model for repeated measures auditory localization data. Applied Statistics, 62(4):515–534, August 2013. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

deCarvalho:2013:DTM

[153] M. de Carvalho, K. F. Turkman, and A. Rua. Dynamic threshold modelling and the US business cycle. Applied Statistics, 62(4):535–550, August 2013. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lunn:2013:FBH

[154] David Lunn, Jessica Barrett, Michael Sweeting, and Simon Thompson. Fully Bayesian hierarchical modelling in two stages, with application to metaanalysis. Applied Statistics, 62(4):551– 572, August 2013. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Melnykov:2013:FMM

[155] Volodymyr Melnykov. Finite mixture modelling in mass spectrometry analysis. Applied Statistics, 62(4):573– 592, August 2013. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lee:2013:LAS

[156] Duncan Lee and Richard Mitchell. Locally adaptive spatial smoothing using conditional auto-regressive models. Applied Statistics, 62(4):593–608, August 2013. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sherlock:2013:CHM

[157] Chris Sherlock, Tatiana Xifara, Sandra Telfer, and Mike Begon. A coupled hidden Markov model for disease interactions. Applied Statistics, 62(4):609– 627, August 2013. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic).

Potter:2013:PLA

[158] Gail E. Potter and Niel Hens. A penalized likelihood approach to estimate within-household contact networks from egocentric data. Applied Statistics, 62 (4):629–648, August 2013. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Corradi:2013:EKI

[159] Fabio Corradi and Federico Ricciardi. Evaluation of kinship identification systems based on short tandem repeat DNA profiles. Applied Statistics, 62(5):649– 668, November 2013. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lawrance:2013:EGF

[160] A. J. Lawrance. Exploratory graphics for financial time series volatility. Applied Statistics, 62(5):669–686, November 2013. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zhang:2013:AHT

[161] Zhiwei Zhang, Chenguang Wang, Lei Nie, and Guoxing Soon. Assessing the heterogeneity of treatment effects via potential outcomes of individual patients. Applied Statistics, 62(5):687–704, November 2013. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Paul:2013:HLM

[162] Sudeshna Paul and A. James O’Malley. Hierarchical longitudinal models of relationships in social networks. Applied Statistics, 62(5):705–722, November 2013. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Noufaily:2013:PQR

[163] Angela Noufaily and M. C. Jones.

Parametric quantile regression based on the generalized gamma distribution. Applied Statistics, 62(5):723–740, November 2013. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Soh:2013:SEL

[164] Patrice Takam Soh, Eugène-Patrice Ndong Nguéma, Henri Gwet, and Michel Ndoumbè-Nkeng. Smooth estimation of a lifetime distribution with competing risks by using regular interval observations: application to cocoa fruits growth. Applied Statistics, 62(5):741– 760, November 2013. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2013:IAI

[165] Anonymous. Index of authors: Index of authors, volume 62, 2013. Applied Statistics, 62(5):761–763, November 2013. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2013:CCV

[166] Anonymous. Contents: Contents of volume 62, 2013. Applied Statistics, 62(5): 765–766, November 2013. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ma:2014:EDO

[167] Yanyuan Ma and Yuanjia Wang. Estimating disease onset distribution functions in mutation carriers with censored mixture data. Applied Statistics, 63(1): 1–23, January 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Guglielmi:2014:SBM

[168] Alessandra Guglielmi, Francesca Ieva, Anna M. Paganoni, Fabrizio Ruggeri, and Jacopo Soriano. Semiparametric Bayesian models for clustering and classification in the presence of unbalanced in-hospital survival. Applied Statistics, 63(1):25–46, January 2014. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

ODonnell:2014:FRM

[169] David O’Donnell, Alastair Rushworth, Adrian W. Bowman, E. Marian Scott, and Mark Hallard. Flexible regression models over river networks. Applied Statistics, 63(1):47–63, January 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Veiga:2014:USW

[170] Alinne Veiga, Peter W. F. Smith, and James J. Brown. The use of sample weights in multivariate multilevel models with an application to income data collected by using a rotating panel survey. Applied Statistics, 63(1):65– 84, January 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Almansa:2014:FMM

[171] Josué Almansa, Jeroen K. Vermunt, Carlos G. Forero, and Jordi Alonso. A factor mixture model for multivariate survival data: an application to the analysis of lifetime mental disorders. Applied Statistics, 63(1):85–102, January 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). European Study of the Epidemiology of Mental Disorders– Mental Health Disability: a European Assessment in the Year 2000 Investigators.

Neto:2014:ASV

[172] Joaquim Henriques Vianna Neto, Alexandra M. Schmidt, and Peter Guttorp. Accounting for spatially varying directional effects in spatial covariance structures. Applied Statistics, 63(1):103– 122, January 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hering:2014:CSC

[173] Amanda S. Hering and Sean Bair. Characterizing spatial and chronological target selection of serial offenders. Applied Statistics, 63(1):123–140, January 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Unkel:2014:TVF

[174] Steffen Unkel, C. Paddy Farrington, Heather J. Whitaker, and Richard Pebody. Time varying frailty models and the estimation of heterogeneities in transmission of infectious diseases. Applied Statistics, 63(1):141–158, January 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cai:2014:BDF

[175] Chunyan Cai, Ying Yuan, and Yuan Ji.

A Bayesian dose finding design for oncology clinical trials of combinational biological agents. Applied Statistics, 63(1): 159–173, January 2014. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Seppa:2014:RVR

[176] Karri Seppä, Timo Hakulinen, and Esa Läärä. Regional variation in relative survival — quantifying the effects of the competing risks of death by using a cure fraction model with random effects. Applied Statistics, 63(1):175–190, January 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kuha:2014:PTI

[177] Jouni Kuha and Irini Moustaki. Preface to the themed issue on social statistics. Applied Statistics, 63(2):191–193, February 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

OMuircheartaigh:2014:GUE

[178] Colm O’Muircheartaigh and Larry V. Hedges. Generalizing from unrepresentative experiments: a stratified propensity score approach. Applied Statistics, 63(2):195–210, February 2014. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Skrondal:2014:HIC

[179] Anders Skrondal and Sophia Rabe-Hesketh. Handling initial conditions and endogenous covariates in dynamic/transition models for binary data with unobserved heterogeneity. Applied Statistics, 63(2):211–237, February 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Washbrook:2014:INI

[180] Elizabeth Washbrook, Paul S. Clarke, and Fiona Steele. Investigating nonignorable dropout in panel studies of residential mobility. Applied Statistics, 63(2):239–266, February 2014. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bartolucci:2014:LAS

[181] Francesco Bartolucci, Silvia Bacci, and Fulvia Pennoni. Longitudinal analysis of self-reported health status by mixture latent auto-regressive models. Applied Statistics, 63(2):267–288, February 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

B:2014:LMM

[182] Jorge González B., Paul De Boeck, and Francis Tuerlinckx. Linear mixed modelling for data from a double mixed factorial design with covariates: a casestudy on semantic categorization response times. Applied Statistics, 63(2): 289–302, February 2014. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Silva:2014:UAI

[183] Damião N. Da Silva and Chris Skinner.

The use of accuracy indicators to correct for survey measurement error. Applied Statistics, 63(2):303–319, February 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kuha:2014:ICM

[184] Jouni Kuha and Jonathan Jackson. The item count method for sensitive survey questions: modelling criminal behaviour. Applied Statistics, 63(2):321– 341, February 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Moustaki:2014:LVM

[185] Irini Moustaki and Martin Knott. Latent variable models that account for atypical responses. Applied Statistics, 63(2):343–360, February 2014. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Slooten:2014:PSF

[186] Klaas Slooten and Ronald Meester. Probabilistic strategies for familial DNA searching. Applied Statistics, 63(3):361– 384, April 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Guindani:2014:BSA

[187] Michele Guindani, Nuno Sepúlveda, Carlos Daniel Paulino, and Peter Müller. A Bayesian semiparametric approach for the differential analysis of sequence counts data. Applied Statistics, 63(3): 385–404, April 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Scheuerer:2014:SAP

[188] Michael Scheuerer and Luca Büermann. Spatially adaptive post-processing of ensemble forecasts for temperature. Applied Statistics, 63(3):405–422, April 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jandarov:2014:EGM

[189] Roman Jandarov, Murali Haran, Ottar Bjørnstad, and Bryan Grenfell. Emulating a gravity model to infer the spatiotemporal dynamics of an infectious disease. Applied Statistics, 63(3):423– 444, April 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Huang:2014:MMC

[190] Xiaobi Huang, Michael R. Elliott, and Siobán D. Harlow. Modelling menstrual cycle length and variability at the approach of menopause by using hierarchical change point models. Applied Statistics, 63(3):445–466, April 2014. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Brown:2014:SIC

[191] Patrick E. Brown, Florencia Chimard, Alexander Remorov, Jeffrey S. Rosenthal, and Xin Wang. Statistical inference and computational efficiency for spatial infectious disease models with plantation data. Applied Statistics, 63(3):467– 482, April 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Feng:2014:GQL

[192] Zeny Feng. A generalized quasilikelihood scoring approach for simultaneously testing the genetic association of multiple traits. Applied Statistics, 63(3): 483–498, April 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dong:2014:OSR

[193] Ting Dong, Liansheng Larry Tang, and William F. Rosenberger. Optimal sampling ratios in comparative diagnostic trials. Applied Statistics, 63(3):499– 514, April 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zhang:2014:SAN

[194] Zhiwei Zhang, Lei Nie, Guoxing Soon, and Bo Zhang. Sensitivity analysis in non-inferiority trials with residual inconstancy after covariate adjustment. Applied Statistics, 63(4):515–538, August 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Savitsky:2014:BNP

[195] Terrance D. Savitsky and Siddhartha R. Dalal. Bayesian non-parametric analysis of multirater ordinal data, with application to prioritizing research goals for prevention of suicide. Applied Statistics, 63 (4):539–557, August 2014. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Xu:2014:TSA

[196] Jiajing Xu and Guosheng Yin. Twostage adaptive randomization for delayed response in clinical trials. Applied Statistics, 63(4):559–578, August 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Erhardt:2014:ISI

[197] Erik Barry Erhardt and Edward J. Bedrick. Inference for stable isotope mixing models: a study of the diet of dunlin. Applied Statistics, 63(4):579– 593, August 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zhang:2014:BHS

[198] Lin Zhang, Veerabhadran Baladandayuthapani, Bani K. Mallick, Ganiraju C. Manyam, Patricia A. Thompson, Melissa L. Bondy, and Kim-Anh Do. Bayesian hierarchical structured variable selection methods with application to molecular inversion probe studies in breast cancer. Applied Statistics, 63 (4):595–620, August 2014. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Follmann:2014:EBP

[199] Dean Follmann, Jing Qin, M. Lourdes Guerrero, J. Gabrielle Breugelmans, Gustave Rosales Pedraza, Bradford D. Gessner, and Guillermo M. Ruiz-Palacios. Estimating the burden of pertussis in Mexican adolescents from paired serological data by using a bivariate mixture model. Applied Statistics, 63 (4):621–637, August 2014. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Balasubramanian:2014:VIM

[200] Raji Balasubramanian, E. Andres Houseman, Brent A. Coull, Michael H. Lev, Lee H. Schwamm, and Rebecca A. Betensky. Variable importance in matched case-control studies in settings of high dimensional data. Applied Statistics, 63(4):639–655, August 2014. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kuk:2014:VTS

[201] Anthony Y. C. Kuk, Jialiang Li, and A. John Rush. Variable and threshold selection to control predictive accuracy in logistic regression. Applied Statistics, 63(4):657–672, August 2014. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Baddeley:2014:MPP

[202] Adrian Baddeley, Aruna Jammalamadaka, and Gopalan Nair. Multitype point process analysis of spines on the dendrite network of a neuron. Applied Statistics, 63(5):673–694, November 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sommer:2014:STT

[203] Julia C. Sommer and Volker J. Schmid. Spatial two-tissue compartment model for dynamic contrast-enhanced magnetic resonance imaging. Applied Statistics, 63(5):695–713, November 2014. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Song:2014:BHD

[204] Yong Song, Yun Li, Bryson Bates, and Christopher K. Wikle. A Bayesian hierarchical downscaling model for southwest Western Australia rainfall. Applied Statistics, 63(5):715–736, November 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Neelon:2014:MSM

[205] Brian Neelon, Alan E. Gelfand, and Marie Lynn Miranda. A multivariate spatial mixture model for areal data: examining regional differences in standardized test scores. Applied Statistics, 63(5): 737–761, November 2014. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nyamundanda:2014:DPP

[206] Gift Nyamundanda, Isobel Claire Gormley, and Lorraine Brennan. A dynamic probabilistic principal components model for the analysis of longitudinal metabolomics data. Applied Statistics, 63(5):763–782, November 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sonksen:2014:ILC

[207] Michael D. Sonksen and Mario Peruggia. Inferences on lung cancer mortality rates based on reference priors under partial ordering. Applied Statistics, 63(5):783–800, November 2014. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2014:IA

[208] Anonymous. Index of authors. Applied Statistics, 63(5):801–804, November 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2014:CCV

[209] Anonymous. Contents of volume. Applied Statistics, 63(5):805–806, November 2014. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Muff:2015:BAM

[210] Stefanie Muff, Andrea Riebler, Leonhard Held, Håvard Rue, and Philippe Saner. Bayesian analysis of measurement error models using integrated nested Laplace approximations. Applied Statistics, 64(2):231–252, February 2015. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lee:2015:BSA

[211] Kyu Ha Lee, Sebastien Haneuse, Deborah Schrag, and Francesca Dominici. Bayesian semiparametric analysis of semicompeting risks data: investigating hospital readmission after a pancreatic cancer diagnosis. Applied Statistics, 64(2):253–273, February 2015. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wilson:2015:EEA

[212] Amy Wilson, Colin Aitken, Richard Sleeman, and James Carter. The evaluation of evidence for auto-correlated data in relation to traces of cocaine on banknotes. Applied Statistics, 64(2):275– 298, February 2015. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Papastathopoulos:2015:SOU

[213] Ioannis Papastathopoulos and Jonathan A. Tawn. Stochastic ordering under conditional modelling of extreme values: drug-induced liver injury. Applied Statistics, 64(2):299–317, February 2015. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cook:2015:GSM

[214] Andrea J. Cook, Robert D. Wellman, Jennifer C. Nelson, Lisa A. Jackson, and Ram C. Tiwari. Group sequential method for observational data by using generalized estimating equations: application to Vaccine Safety Datalink. Applied Statistics, 64(2):319–338, February 2015. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

McLain:2015:JAL

[215] Alexander C. McLain, Rajeshwari Sundaram, and Germaine M. Buck Louis. Joint analysis of longitudinal and survival data measured on nested timescales by using shared parameter models: an application to fecundity data. Applied Statistics, 64(2):339– 357, February 2015. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lynch:2015:PTT

[216] Miranda L. Lynch and Victor DeGruttola. Predicting time to threshold for initiating antiretroviral treatment to evaluate cost of treatment as prevention of human immunodeficiency virus. Applied Statistics, 64(2):359–375, February 2015. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Keim:2015:ACS

[217] Michael J. Keim and Donald B. Percival. Assessing characteristic scales using wavelets. Applied Statistics, 64(2):377– 393, February 2015. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Obermeier:2015:FDL

[218] Viola Obermeier, Fabian Scheipl, Christian Heumann, Joachim Wassermann, and Helmut Küchenhoff. Flexible distributed lags for modelling earthquake data. Applied Statistics, 64(2):395– 412, February 2015. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kim:2015:JTO

[219] Sungwook Kim and Nancy Flournoy.

Optimal experimental design for systems with bivariate failures under a bivariate Weibull function. Applied Statistics, 64(3):413–432, April 2015. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tang:2015:JTO

[220] Li Tang, Robert H. Lyles, Caroline C.

King, Joseph W. Hogan, and Yungtai Lo. Regression analysis for differentially misclassified correlated binary outcomes. Applied Statistics, 64(3):433– 449, April 2015. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mehtala:2015:OOT

[221] Juha Mehtälä, Kari Auranen, and Sangita Kulathinal. Optimal observation times for multistate Markov models — applications to pneumococcal colonization studies. Applied Statistics, 64(3): 451–468, April 2015. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chen:2015:JTO

[222] Yong Chen, Yulun Liu, Jing Ning, Janice Cormier, and Haitao Chu. A hybrid model for combining case–control and cohort studies in systematic reviews of diagnostic tests. Applied Statistics, 64 (3):469–489, April 2015. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Haggarty:2015:JTO

[223] R. A. Haggarty, C. A. Miller, and E. M. Scott. Spatially weighted functional clustering of river network data. Applied Statistics, 64(3):491–506, April 2015. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Liu:2015:JTO

[224] Suyu Liu and Ying Yuan. Bayesian optimal interval designs for phase I clinical trials. Applied Statistics, 64(3):507– 523, April 2015. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Leclerc:2015:JTO

[225] Martin Leclerc, Embrace Investigators, Gemo Study Collaborators, Inherit Investigators, Antonis C. Antoniou, Jacques Simard, and Lajmi Lakhal-Chaieb. Analysis of multivariate failure times in the presence of selection bias with application to breast cancer. Applied Statistics, 64(3):525–541, April 2015. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Heaps:2015:JTO

[226] Sarah E. Heaps, Richard J. Boys, and Malcolm Farrow. Bayesian modelling of rainfall data by using non-homogeneous hidden Markov models and latent Gaussian variables. Applied Statistics, 64(3): 543–568, April 2015. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Klein:2015:BSA

[227] Nadja Klein, Thomas Kneib, Stephan Klasen, and Stefan Lang. Bayesian structured additive distributional regression for multivariate responses. Applied Statistics, 64(4):569–591, August 2015. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Swartz:2015:BAA

[228] Tim B. Swartz, Paramjit S. Gill, and Saman Muthukumarana. A Bayesian approach for the analysis of triadic data in cognitive social structures. Applied Statistics, 64(4):593–610, August 2015. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sewell:2015:AFS

[229] Daniel K. Sewell and Yuguo Chen. Analysis of the formation of the structure of social networks by using latent space models for ranked dynamic networks. Applied Statistics, 64(4):611– 633, August 2015. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zang:2015:OMA

[230] Yong Zang, Suyu Liu, and Ying Yuan. Optimal marker-adaptive designs for targeted therapy based on imperfectly measured biomarkers. Applied Statistics, 64(4):635–650, August 2015. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lavigne:2015:AEC

[231] A. Lavigne, N. Eckert, L. Bel, and E. Parent. Adding expert contributions to the spatiotemporal modelling of avalanche activity under different climatic influences. Applied Statistics, 64 (4):651–671, August 2015. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Schnell:2015:MCR

[232] Patrick Schnell, Dipankar Bandyopadhyay, Brian J. Reich, and Martha Nunn. A marginal cure rate proportional hazards model for spatial survival data. Applied Statistics, 64(4):673–691, August 2015. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Black:2015:ORC

[233] Michael S. Black, Christopher R. Bilder, and Joshua M. Tebbs. Optimal retesting configurations for hierarchical group testing. Applied Statistics, 64(4):693– 710, August 2015. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shih:2015:MTT

[234] Joanna H. Shih, Paul S. Albert, Pauline Mendola, and Katherine L. Grantz. Modelling the type and timing of consecutive events: application to predicting preterm birth in repeated pregnancies. Applied Statistics, 64(5):711–730, November 2015. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jiang:2015:MSL

[235] Bei Jiang, Naisyin Wang, Mary D. Sammel, and Michael R. Elliott. Modelling short- and long-term characteristics of follicle stimulating hormone as predictors of severe hot flashes in the Penn Ovarian Aging Study. Applied Statistics, 64(5):731–753, November 2015. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bakar:2015:SMA

[236] K. Shuvo Bakar, Philip Kokic, and Huidong Jin. A spatiodynamic model for assessing frost risk in south-eastern Australia. Applied Statistics, 64(5):755–778, November 2015. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rodriguez:2015:MAS

[237] Abel Rodríguez and Scott Moser. Measuring and accounting for strategic abstentions in the US Senate, 1989– 2012. Applied Statistics, 64(5):779–797, November 2015. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Quick:2015:HCA

[238] Harrison Quick, Bradley P. Carlin, and Sudipto Banerjee. Heteroscedastic conditional auto-regression models for areally referenced temporal processes for analysing California asthma hospitalization data. Applied Statistics, 64(5): 799–813, November 2015. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Long:2015:MZI

[239] D. Leann Long, John S. Preisser, Amy H. Herring, and Carol E. Golin. A marginalized zero-inflated Poisson regression model with random effects. Applied Statistics, 64(5):815–830, November 2015. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2015:IAV

[240] Anonymous. Index of authors, volume 64, 2015. Applied Statistics, 64(5):831– 834, November 2015. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2015:CV

[241] Anonymous. Contents of volume 64, 2015. Applied Statistics, 64(5):835–836, November 2015. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kazianka:2016:OBE

[242] Hannes Kazianka. Objective Bayesian estimation of the probability of default. Applied Statistics, 65(1):1–27, January 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sykulski:2016:LTS

[243] Adam M. Sykulski, Sofia C. Olhede, Jonathan M. Lilly, and Eric Danioux. Lagrangian time series models for ocean surface drifter trajectories. Applied Statistics, 65(1):29–50, January 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Geraci:2016:PPC

[244] Marco Geraci and Alessio Farcomeni.

Probabilistic principal component analysis to identify profiles of physical activity behaviours in the presence of nonignorable missing data. Applied Statistics, 65(1):51–75, January 2016. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Waagepetersen:2016:AMP

[245] Rasmus Waagepetersen, Yongtao Guan, Abdollah Jalilian, and Jorge Mateu. Analysis of multispecies point patterns by using multivariate log-Gaussian Cox processes. Applied Statistics, 65(1):77– 96, January 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wu:2016:PLC

[246] Zhenke Wu, Maria Deloria-Knoll, Laura L. Hammitt, Scott L. Zeger, and for the Pneumonia Etiology Research for Child Health Core Team. Partially latent class models for case-control studies of childhood pneumonia aetiology. Applied Statistics, 65(1):97–114, January 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zhu:2016:ECD

[247] Yeying Zhu, Debashis Ghosh, Donna L.

Coffman, and Jennifer S. Savage. Estimating controlled direct effects of restrictive feeding practices in the ‘Early dieting in girls’ study. Applied Statistics, 65(1):115–130, January 2016. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Josefsson:2016:CIL

[248] Maria Josefsson, Xavier de Luna, Michael J. Daniels, and Lars Nyberg. Causal inference with longitudinal outcomes and non-ignorable dropout: estimating the effect of living alone on cognitive decline. Applied Statistics, 65(1): 131–144, January 2016. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

White:2016:PHP

[249] Nicole White and Kerrie Mengersen. Predicting health programme participation: a gravity-based, hierarchical modelling approach. Applied Statistics, 65(1):145–166, January 2016. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

vandenHout:2016:JMD

[250] Ardo van den Hout and Graciela Muniz-Terrera. Joint models for discrete longitudinal outcomes in aging research. Applied Statistics, 65(1):167–186, January 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zhang:2016:BSM

[251] Ran Zhang, Claudia Czado, and Karin Sigloch. Bayesian spatial modelling for high dimensional seismic inverse problems. Applied Statistics, 65(2):187– 213, February 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Goldsmith:2016:ASE

[252] Jeff Goldsmith and Tomoko Kitago. Assessing systematic effects of stroke on motor control by using hierarchical function-on-scalar regression. Applied Statistics, 65(2):215–236, February 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dureau:2016:BAE

[253] J. Dureau, K. Kalogeropoulos, P. Vickerman, M. Pickles, and M.-C. Boily. A Bayesian approach to estimate changes in condom use from limited human immunodeficiency virus prevalence data. Applied Statistics, 65(2):237–257, February 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Guo:2016:DSF

[254] Beibei Guo, Yisheng Li, and Ying Yuan.

A dose-schedule finding design for phase I–II clinical trials. Applied Statistics, 65(2):259–272, February 2016. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hobbs:2016:BGS

[255] Brian P. Hobbs, Peter F. Thall, and Steven H. Lin. Bayesian group sequential clinical trial design using total toxicity burden and progression-free survival. Applied Statistics, 65(2):273– 297, February 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Winkelmann:2016:FGP

[256] Lars Winkelmann. Forward guidance and the predictability of monetary policy: a wavelet-based jump detection approach. Applied Statistics, 65(2):299– 314, February 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Steiner:2016:CTB

[257] Stefan H. Steiner, Oana Danila, Cecilia A. Cotton, Daniel Severn, and R. Jock Mackay. Comparing two binary diagnostic tests with repeated measurements. Applied Statistics, 65(2):315– 329, February 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Misumi:2016:MEH

[258] Toshihiro Misumi and Sadanori Konishi. Mixed effects historical varyingcoefficient model for evaluating doseresponse in flexible dose trials. Applied Statistics, 65(2):331–344, February 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Winter:2016:MHC

[259] Hugo C. Winter and Jonathan A. Tawn. Modelling heatwaves in central France: a case-study in extremal dependence. Applied Statistics, 65(3):345–365, April 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Heydari:2016:BHM

[260] Jonathan Heydari, Conor Lawless, David A. Lydall, and Darren J. Wilkinson. Bayesian hierarchical modelling for inferring genetic interactions in yeast. Applied Statistics, 65(3):367–393, April 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pomann:2016:TSD

[261] Gina-Maria Pomann, Ana-Maria Staicu, and Sujit Ghosh. A two-sample distribution-free test for functional data with application to a diffusion tensor imaging study of multiple sclerosis.

Applied Statistics, 65(3):395–414, April 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Schepsmeier:2016:DMR

[262] Ulf Schepsmeier and Claudia Czado. Dependence modelling with regular vine copula models: a case-study for car crash simulation data. Applied Statistics, 65(3):415–429, April 2016. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Yang:2016:TSM

[263] Hanyu Yang, Runze Li, Robert A.

Zucker, and Anne Buu. Two-stage model for time varying effects of zeroinflated count longitudinal covariates with applications in health behaviour research. Applied Statistics, 65(3):431– 444, April 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rivest:2016:GAR

[264] Louis-Paul Rivest, Thierry Duchesne, Aurélien Nicosia, and Daniel Fortin. A general angular regression model for the analysis of data on animal movement in ecology. Applied Statistics, 65(3):445– 463, April 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Xu:2016:BTS

[265] Jiajing Xu, Guosheng Yin, David Ohlssen, and Frank Bretz. Bayesian twostage dose finding for cytostatic agents via model adaptation. Applied Statistics, 65(3):465–482, April 2016. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Overstall:2016:MEC

[266] Antony M. Overstall and David C. Woods. Multivariate emulation of computer simulators: model selection and diagnostics with application to a humanitarian relief model. Applied Statistics, 65(4):483–505, August 2016. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kerm:2016:DQW

[267] Philippe Van Kerm, Seunghee Yu, and Chung Choe. Decomposing quantile wage gaps: a conditional likelihood approach. Applied Statistics, 65(4):507– 527, August 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hoegh:2016:SMF

[268] Andrew Hoegh, Marco A. R. Ferreira, and Scotland Leman. Spatiotemporal model fusion: multiscale modelling of civil unrest. Applied Statistics, 65(4): 529–545, August 2016. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lee:2016:BII

[269] Juhee Lee, Peter Müller, Subhajit Sengupta, Kamalakar Gulukota, and Yuan Ji. Bayesian inference for intratumour heterogeneity in mutations and copy number variation. Applied Statistics, 65 (4):547–563, August 2016. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bothmann:2016:RCF

[270] Ludwig Bothmann, Michael Windmann, and Göran Kauermann. Realtime classification of fish in underwater sonar videos. Applied Statistics, 65(4):565– 584, August 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zang:2016:TSM

[271] Yong Zang, J. Jack Lee, and Ying Yuan.

Two-stage marker-stratified clinical trial design in the presence of biomarker misclassification. Applied Statistics, 65(4): 585–601, August 2016. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Baker:2016:CWE

[272] Rose Baker, David Forrest, and Levi Pérez. The compatriot win effect on national sales of a multicountry lottery. Applied Statistics, 65(4):603– 618, August 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bandyopadhyay:2016:NPS

[273] Dipankar Bandyopadhyay and Antonio Canale. Non-parametric spatial models for clustered ordered periodontal data. Applied Statistics, 65(4):619– 640, August 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Liu:2016:CWK

[274] W. Liu, Y. Han, F. Bretz, F. Wan, and P. Yang. Counting by weighing: know your numbers with confidence. Applied Statistics, 65(4):641–648, August 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bonat:2016:MCG

[275] Wagner Hugo Bonat and Bent Jørgensen.

Multivariate covariance generalized linear models. Applied Statistics, 65(5): 649–675, November 2016. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mak:2016:RCP

[276] Simon Mak, Derek Bingham, and Yi Lu. A regional compound Poisson process for hurricane and tropical storm damage. Applied Statistics, 65(5):677–703, November 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sosina:2016:SGG

[277] Sobambo Sosina, Tirthankar Dasgupta, and Qiang Huang. A stochastic graphene growth kinetics model. Applied Statistics, 65(5):705–729, November 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bergen:2016:MME

[278] Silas Bergen, Lianne Sheppard, Joel D. Kaufman, and Adam A. Szpiro. Multipollutant measurement error in air pollution epidemiology studies arising from predicting exposures with penalized regression splines. Applied Statistics, 65 (5):731–753, November 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Li:2016:QRL

[279] Ruosha Li, Xuelin Huang, and Jorge Cortes. Quantile residual life regression with longitudinal biomarker measurements for dynamic prediction. Applied Statistics, 65(5):755–773, November 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Doretti:2016:TNI

[280] Marco Doretti, Sara Geneletti, and Elena Stanghellini. Tackling nonignorable dropout in the presence of time varying confounding. Applied Statistics, 65(5):775–795, November 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Barrett:2016:IAC

[281] James E. Barrett. Information-adaptive clinical trials: a selective recruitment design. Applied Statistics, 65(5):797–808, November 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2016:IAV

[282] Anonymous. Index of authors, volume 65, 2016. Applied Statistics, 65(5):809– 811, November 2016. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2016:CV

[283] Anonymous. Contents of volume 65, 2016. Applied Statistics, 65(5):813–814, November 2016. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2017:IIa

[284] Anonymous. Issue information. Applied Statistics, 66(1):1–2, January 2017. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jandarov:2017:NPC

[285] Roman A. Jandarov, Lianne A. Sheppard, Paul D. Sampson, and Adam A. Szpiro. A novel principal component analysis for spatially misaligned multivariate air pollution data. Applied Statistics, 66(1):3–28, January 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lerch:2017:SBS

[286] Sebastian Lerch and Sándor Baran. Similarity-based semilocal estimation of post-processing models. Applied Statistics, 66(1):29–51, January 2017. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dette:2017:ODT

[287] Holger Dette, Laura Hoyden, Sonja Kuhnt, and Kirsten Schorning. Optimal designs for thermal spraying. Applied Statistics, 66(1):53–72, January 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Saleh:2017:DOE

[288] Moein Saleh, Ming-Hung Kao, and Rong Pan. Design D-optimal event-related functional magnetic resonance imaging experiments. Applied Statistics, 66(1): 73–91, January 2017. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Northrop:2017:CVE

[289] Paul J. Northrop, Nicolas Attalides, and Philip Jonathan. Cross-validatory extreme value threshold selection and uncertainty with application to ocean storm severity. Applied Statistics, 66 (1):93–120, January 2017. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Foreman:2017:SAF

[290] Kyle J. Foreman, Guangquan Li, Nicky Best, and Majid Ezzati. Small area forecasts of cause-specific mortality: application of a Bayesian hierarchical model to US vital registration data. Applied Statistics, 66(1):121–139, January 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rushworth:2017:ASS

[291] Alastair Rushworth, Duncan Lee, and Christophe Sarran. An adaptive spatiotemporal smoothing model for estimating trends and step changes in disease risk. Applied Statistics, 66(1):141– 157, January 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Choi:2017:ECE

[292] Jaeun Choi and A. James O’Malley. Estimating the causal effect of treatment in observational studies with survival time end points and unmeasured confounding. Applied Statistics, 66(1):159– 185, January 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kaizer:2017:IOA

[293] Alexander M. Kaizer and Joseph S.

Koopmeiners. Identifying optimal approaches to early termination in twostage biomarker validation studies. Applied Statistics, 66(1):187–199, January 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Thall:2017:PDS

[294] Peter F. Thall, Hoang Q. Nguyen, and Ralph G. Zinner. Parametric dose standardization for optimizing two-agent combinations in a phase I–II trial with ordinal outcomes. Applied Statistics, 66 (1):201–224, January 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2017:IIb

[295] Anonymous. Issue information. Applied Statistics, 66(2):225–226, February 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lupparelli:2017:LML

[296] Monia Lupparelli and Alberto Roverato. Log-mean linear regression models for binary responses with an application to multimorbidity. Applied Statistics, 66(2):227–252, February 2017. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gouveia:2017:BAR

[297] Sónia Gouveia, Manuel G. Scotto, Christian H. Weiß, and Paulo Jorge S. G. Ferreira. Binary auto-regressive geometric modelling in a DNA context. Applied Statistics, 66(2):253–271, February 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dodd:2017:PSD

[298] Erengul Dodd and George Streftaris. Prediction of settlement delay in critical illness insurance claims by using the generalized beta of the second kind distribution. Applied Statistics, 66(2):273– 294, February 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ritz:2017:SIM

[299] Christian Ritz, Rikke Pilmann Laursen, and Camilla Trab Damsgaard. Simultaneous inference for multilevel linear mixed models — with an application to a large-scale school meal study. Applied Statistics, 66(2):295–311, February 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Atem:2017:LRR

[300] Folefac D. Atem, Jing Qian, Jacqueline E. Maye, Keith A. Johnson, and Rebecca A. Betensky. Linear regression with a randomly censored covariate: application to an Alzheimer’s study. Applied Statistics, 66(2):313–328, February 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Castruccio:2017:ESA

[301] Stefano Castruccio and Joseph Guinness. An evolutionary spectrum approach to incorporate large-scale geographical descriptors on global processes. Applied Statistics, 66(2):329– 344, February 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zhang:2017:SSA

[302] Zhiwei Zhang, Meijuan Li, Min Lin, Guoxing Soon, Tom Greene, and Changyu Shen. Subgroup selection in adaptive signature designs of confirmatory clinical trials. Applied Statistics, 66(2):345–361, February 2017. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ruseckaite:2017:BOC

[303] Aiste Ruseckaite, Peter Goos, and Dennis Fok. Bayesian D-optimal choice designs for mixtures. Applied Statistics, 66(2):363–386, February 2017. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mayrink:2017:BHM

[304] Vinícius Diniz Mayrink and Flávio Bambirra Gonçalves. A Bayesian hidden Markov mixture model to detect overexpressed chromosome regions. Applied Statistics, 66(2):387–412, February 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pak:2017:MMC

[305] Daewoo Pak, Chenxi Li, David Todem, and Woosung Sohn. A multistate model for correlated interval-censored life history data in caries research. Applied Statistics, 66(2):413–423, February 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Godwin:2017:EPS

[306] Ryan T. Godwin and Dankmar Böhning. Estimation of the population size by using the one-inflated positive Poisson model. Applied Statistics, 66(2):425– 448, February 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2017:IIc

[307] Anonymous. Issue information. Applied Statistics, 66(3):449–450, April 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Vinciotti:2017:PTI

[308] Veronica Vinciotti and Ernst Wit. Preface to the themed issue on ‘Statistical network science and its applications’. Applied Statistics, 66(3):451–453, April 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Parker:2017:ODE

[309] Ben M. Parker, Steven G. Gilmour, and John Schormans. Optimal design of experiments on connected units with application to social networks. Applied Statistics, 66(3):455–480, April 2017. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Karwa:2017:SSN

[310] Vishesh Karwa, Pavel N. Krivitsky, and Aleksandra B. Slavković. Sharing social network data: differentially private estimation of exponential family randomgraph models. Applied Statistics, 66(3): 481–500, April 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gile:2017:ANM

[311] Krista J. Gile and Mark S. Handcock.

Analysis of networks with missing data with application to the National Longitudinal Study of Adolescent Health. Applied Statistics, 66(3):501–519, April 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Manitz:2017:SEP

[312] Juliane Manitz, Jonas Harbering, Marie Schmidt, Thomas Kneib, and Anita Schöbel. Source estimation for propagation processes on complex networks with an application to delays in public transportation systems. Applied Statistics, 66 (3):521–536, April 2017. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Arpino:2017:IPS

[313] Bruno Arpino, Luca De Benedictis, and Alessandra Mattei. Implementing propensity score matching with network data: the effect of the General Agreement on Tariffs and Trade on bilateral trade. Applied Statistics, 66(3):537– 554, April 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Durante:2017:BMN

[314] Daniele Durante, Sally Paganin, Bruno Scarpa, and David B. Dunson. Bayesian modelling of networks in complex business intelligence problems. Applied Statistics, 66(3):555–580, April 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Barigozzi:2017:NAV

[315] Matteo Barigozzi and Marc Hallin. A network analysis of the volatility of high dimensional financial series. Applied Statistics, 66(3):581–605, April 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cribben:2017:EWB

[316] Ivor Cribben and Yi Yu. Estimating whole-brain dynamics by using spectral clustering. Applied Statistics, 66(3):607– 627, April 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mohammadi:2017:BMD

[317] Abdolreza Mohammadi, Fentaw Abegaz, Edwin van den Heuvel, and Ernst C. Wit. Bayesian modelling of Dupuytren disease by using Gaussian copula graphical models. Applied Statistics, 66(3): 629–645, April 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Roverato:2017:NPC

[318] Alberto Roverato and Robert Castelo.

The networked partial correlation and its application to the analysis of genetic interactions. Applied Statistics, 66(3): 647–665, April 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2017:IId

[319] Anonymous. Issue information. Applied Statistics, 66(4):667–668, August 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Yiu:2017:EES

[320] Sean Yiu, Vernon T. Farewell, and Brian D. M. Tom. Exploring the existence of a stayer population with mover–stayer counting process models: application to joint damage in psoriatic arthritis. Applied Statistics, 66(4):669–690, August 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wang:2017:HMM

[321] Ting Wang, Jiancang Zhuang, Kazushige Obara, and Hiroshi Tsuruoka. Hidden Markov modelling of sparse time series from non-volcanic tremor observations. Applied Statistics, 66(4):691– 715, August 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Andrianakis:2017:HMC

[322] I. Andrianakis, I. Vernon, N. McCreesh, T. J. McKinley, J. E. Oakley, R. N. Nsubuga, M. Goldstein, and R. G. White. History matching of a complex epidemiological model of human immunodeficiency virus transmission by using variance emulation. Applied Statistics, 66(4):717–740, August 2017. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wang:2017:CMA

[323] Wei Wang and Jeffrey M. Albert. Causal mediation analysis for the Cox proportional hazards model with a smooth baseline hazard estimator. Applied Statistics, 66(4):741–757, August 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mercatanti:2017:DDC

[324] Andrea Mercatanti and Fan Li. Do debit cards decrease cash demand?: causal inference and sensitivity analysis using principal stratification. Applied Statistics, 66(4):759–776, August 2017. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Saint-Cyr:2017:MSF

[325] Legrand D. F. Saint-Cyr and Laurent Piet. Movers and stayers in the farming sector: accounting for unobserved heterogeneity in structural change. Applied Statistics, 66(4):777–795, August 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chen:2017:BCT

[326] Cathy W. S. Chen and Sangyeol Lee. Bayesian causality test for integervalued time series models with applications to climate and crime data. Applied Statistics, 66(4):797–814, August 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

McCrea:2017:NSD

[327] Rachel S. McCrea, Byron J. T. Morgan, and Olivier Gimenez. A new strategy for diagnostic model assessment in capture– recapture. Applied Statistics, 66(4):815– 831, August 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jiang:2017:SOS

[328] Fei Jiang, Yanyuan Ma, and J. Jack Lee. A second-order semiparametric method for survival analysis, with application to an acquired immune deficiency syndrome clinical trial study. Applied Statistics, 66(4):833–846, August 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ma:2017:BDC

[329] Tianzhou Ma, Faming Liang, and George C. Tseng. Biomarker detection and categorization in ribonucleic acid sequencing meta-analysis using Bayesian hierarchical models. Applied Statistics, 66(4):847–867, August 2017. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ferguson:2017:SIM

[330] Elaine A. Ferguson, Jason Matthiopoulos, Robert H. Insall, and Dirk Husmeier. Statistical inference of the mechanisms driving collective cell movement. Applied Statistics, 66(4):869– 890, August 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2017:II

[331] Anonymous. Issue information. Applied Statistics, 66(5):891–892, November 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Camarda:2017:MTD

[332] Carlo G. Camarda, Paul H. C. Eilers, and Jutta Gampe. Modelling trends in digit preference patterns. Applied Statistics, 66(5):893–918, November 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cunha:2017:NSS

[333] Marcelo Cunha, Dani Gamerman, Montserrat Fuentes, and Marina Paez. A non-stationary spatial model for temperature interpolation applied to the state of Rio de Janeiro. Applied Statistics, 66(5):919–939, November 2017. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jalbert:2017:SME

[334] Jonathan Jalbert, Anne-Catherine Favre, Claude Bélisle, and Jean-François Angers. A spatiotemporal model for extreme precipitation simulated by a climate model, with an application to assessing changes in return levels over North America. Applied Statistics, 66 (5):941–962, November 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Willis:2017:IDC

[335] Amy Willis, John Bunge, and Thea Whitman. Improved detection of changes in species richness in high diversity microbial communities. Applied Statistics, 66(5):963–977, November 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pan:2017:CPP

[336] Haitao Pan, Ying Yuan, and Jielai Xia.

A calibrated power prior approach to borrow information from historical data with application to biosimilar clinical trials. Applied Statistics, 66(5):979–996, November 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hole:2017:UHO

[337] Arne Risa Hole and Hong Il Yoo. The use of heuristic optimization algorithms to facilitate maximum simulated likelihood estimation of random parameter logit models. Applied Statistics, 66(5): 997–1013, November 2017. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Iasonos:2017:PDA

[338] Alexia Iasonos and John O’Quigley.

Phase I designs that allow for uncertainty in the attribution of adverse events. Applied Statistics, 66(5):1015– 1030, November 2017. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Foster:2017:EGS

[339] Scott D. Foster, Nicole A. Hill, and Mitchell Lyons. Ecological grouping of survey sites when sampling artefacts are present. Applied Statistics, 66(5):1031– 1047, November 2017. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ivanova:2017:MMD

[340] Anna Ivanova, Geert Molenberghs, and Geert Verbeke. Mechanism for missing data incorporated in joint modelling of ordinal responses. Applied Statistics, 66(5):1049–1064, November 2017. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

DeStavola:2017:DBA

[341] Bianca L. De Stavola and D. R. Cox. Detecting bias arising from delayed recording of time. Applied Statistics, 66 (5):1065–1073, November 2017. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Yang:2017:JRA

[342] Chi-Chuan Yang, Yi-Hau Chen, and Hsing-Yi Chang. Joint regression analysis of marginal quantile and quantile association: application to longitudinal body mass index in adolescents. Applied Statistics, 66(5):1075–1090, November 2017. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2017:IAV

[343] Anonymous. Index of authors, volume 66, 2017. Applied Statistics, 66(5):1091– 1094, November 2017. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2017:CV

[344] Anonymous. Contents of volume 66, 2017. Applied Statistics, 66(5):1095– 1097, November 2017. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2018:IIa

[345] Anonymous. Issue information. Applied Statistics, 67(1):1–2, January 2018. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Papastamoulis:2018:BMS

[346] Panagiotis Papastamoulis and Magnus Rattray. A Bayesian model selection approach for identifying differentially expressed transcripts from RNA sequencing data. Applied Statistics, 67(1):3– 23, January 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Carson:2018:BMS

[347] Jake Carson, Michel Crucifix, Simon Preston, and Richard D. Wilkinson. Bayesian model selection for the glacial– interglacial cycle. Applied Statistics, 67 (1):25–54, January 2018. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pini:2018:DSF

[348] Alessia Pini, Simone Vantini, Bianca Maria Colosimo, and Marco Grasso. Domainselective functional analysis of variance for supervised statistical profile monitoring of signal data. Applied Statistics, 67 (1):55–81, January 2018. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mortensen:2018:UHR

[349] Jacob W. Mortensen, Matthew J.

Heaton, and Olga V. Wilhelmi. Urban heat risk mapping using multiple point patterns in Houston, Texas. Applied Statistics, 67(1):83–102, January 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gran:2018:ETE

[350] Jon Michael Gran, Rune Hoff, Kjetil Røysland, Bruno Ledergerber, James Young, and Odd O. Aalen. Estimating the treatment effect on the treated under time-dependent confounding in an application to the Swiss HIV Cohort Study. Applied Statistics, 67(1):103– 125, January 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Liu:2018:BMT

[351] Yulun Liu, Stacia M. DeSantis, and Yong Chen. Bayesian mixed treatment comparisons meta-analysis for correlated outcomes subject to reporting bias. Applied Statistics, 67(1):127– 144, January 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Li:2018:AID

[352] Qiuju Li and Li Su. Accommodating informative dropout and death: a joint modelling approach for longitudinal and semicompeting risks data. Applied Statistics, 67(1):145–163, January 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stathopoulos:2018:BEC

[353] Vassilios Stathopoulos, Veronica Zamora-Gutierrez, Kate E. Jones, and Mark Girolami. Bat echolocation call identification for biodiversity monitoring: a probabilistic approach. Applied Statistics, 67(1):165–183, January 2018. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hasan:2018:SMC

[354] M. Tariqul Hasan, Gary Sneddon, and Renjun Ma. Simultaneously modelling clustered marginal counts and multinomial proportions with zero inflation with application to analysis of osteoporotic fractures data. Applied Statistics, 67(1): 185–200, January 2018. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kashlak:2018:MMO

[355] Adam B. Kashlak, Eoin Devane, Helge Dietert, and Henry Jackson. Markov models for ocular fixation locations in the presence and absence of colour. Applied Statistics, 67(1):201–215, January 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Farcomeni:2018:FGC

[356] Alessio Farcomeni. Fully general Chao and Zelterman estimators with application to a whale shark population. Applied Statistics, 67(1):217–229, January 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shaddick:2018:DIM

[357] Gavin Shaddick, Matthew L. Thomas, Amelia Green, Michael Brauer, Aaron van Donkelaar, Rick Burnett, Howard H. Chang, Aaron Cohen, Rita Van Dingenen, Carlos Dora, Sophie Gumy, Yang Liu, Randall Martin, Lance A. Waller, Jason West, James V. Zidek, and Annette Prüss-Ustün. Data integration model for air quality: a hierarchical approach to the global estimation of exposures to ambient air pollution. Applied Statistics, 67(1):231–253, January 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chaurasia:2018:PMM

[358] Ashok Chaurasia, Danping Liu, and Paul S. Albert. Pattern–mixture models with incomplete informative cluster size: application to a repeated pregnancy study. Applied Statistics, 67(1): 255–273, January 2018. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Misumi:2018:SEB

[359] Munechika Misumi, Kyoji Furukawa, John B. Cologne, and Harry M. Cullings. Simulation–extrapolation for bias correction with exposure uncertainty in radiation risk analysis utilizing grouped data. Applied Statistics, 67(1): 275–289, January 2018. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Quick:2018:MST

[360] Harrison Quick, Lance A. Waller, and Michele Casper. A multivariate space– time model for analysing county level heart disease death rates by race and sex. Applied Statistics, 67(1):291–304, January 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2018:IIb

[361] Anonymous. Issue information. Applied Statistics, 67(2):305–306, February 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Liao:2018:SAF

[362] Xiaomei Liao, Xin Zhou, Molin Wang, Jaime E. Hart, Francine Laden, and Donna Spiegelman. Survival analysis with functions of mismeasured covariate histories: the case of chronic air pollution exposure in relation to mortality in the nurses’ health study. Applied Statistics, 67(2):307–327, February 2018. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Adegboye:2018:ASD

[363] Oyelola A. Adegboye, Denis H. Y. Leung, and You-Gan Wang. Analysis of spatial data with a nested correlation structure. Applied Statistics, 67(2):329– 354, February 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Signorelli:2018:PIA

[364] Mirko Signorelli and Ernst C. Wit. A penalized inference approach to stochastic block modelling of community structure in the Italian Parliament. Applied Statistics, 67(2):355–369, February 2018. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Guo:2018:MAC

[365] Zijian Guo, Dylan S. Small, Stuart A.

Gansky, and Jing Cheng. Mediation analysis for count and zero-inflated count data without sequential ignorability and its application in dental studies. Applied Statistics, 67(2):371–394, February 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ouni:2018:CRP

[366] Z. Ouni, C. Denis, C. Chauvel, and A. Chambaz. Contextual ranking by passive safety of generational classes of light vehicles. Applied Statistics, 67(2): 395–416, February 2018. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wu:2018:DBN

[367] Paul P.-Y. Wu, M. Julian Caley, Gary A. Kendrick, Kathryn McMahon, and Kerrie Mengersen. Dynamic Bayesian network inferencing for non-homogeneous complex systems. Applied Statistics, 67(2):417–434, February 2018. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wilms:2018:MVA

[368] Ines Wilms, Luca Barbaglia, and Christophe Croux. Multiclass vector auto-regressive models for multistore sales data. Applied Statistics, 67(2):435– 452, February 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

deSouza:2018:GAM

[369] Juliana B. de Souza, Valdério A. Reisen, Glaura C. Franco, Márton Ispány, Pascal Bondon, and Jane Meri Santos. Generalized additive models with principal component analysis: an application to time series of respiratory disease and air pollution data. Applied Statistics, 67(2): 453–480, February 2018. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Yiu:2018:CMM

[370] Sean Yiu, Vernon T. Farewell, and Brian D. M. Tom. Clustered multistate models with observation level random effects, mover–stayer effects and dynamic covariates: modelling transition intensities and sojourn times in a study of psoriatic arthritis. Applied Statistics, 67(2):481– 500, February 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beltran:2018:EDM

[371] Daniel O. Beltran and David Draper.

Estimating dynamic macroeconomic models: how informative are the data? Applied Statistics, 67(2):501–520, February 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2018:IIc

[372] Anonymous. Issue information. Applied Statistics, 67(3):521–522, April 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Valle:2018:BNP

[373] Luciana Dalla Valle, Fabrizio Leisen, and Luca Rossini. Bayesian nonparametric conditional copula estimation of twin data. Applied Statistics, 67 (3):523–548, April 2018. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gallo:2018:CSS

[374] Giampiero M. Gallo and Edoardo Otranto. Combining sharp and smooth transitions in volatility dynamics: a fuzzy regime approach. Applied Statistics, 67(3):549–573, April 2018. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ranalli:2018:SSC

[375] Monia Ranalli, Francesco Lagona, Marco Picone, and Enrico Zambianchi. Segmentation of sea current fields by cylindrical hidden Markov models: a composite likelihood approach. Applied Statistics, 67(3):575–598, April 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gloaguen:2018:SDE

[376] Pierre Gloaguen, Marie-Pierre Etienne, and Sylvain Le Corff. Stochastic differential equation based on a multimodal potential to model movement data in ecology. Applied Statistics, 67(3):599– 619, April 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rugamer:2018:BFS

[377] David Rügamer, Sarah Brockhaus, Kornelia Gentsch, Klaus Scherer, and Sonja Greven. Boosting factor-specific functional historical models for the detection of synchronization in bioelectrical signals. Applied Statistics, 67(3):621– 642, April 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Erhardt:2018:SDI

[378] Tobias M. Erhardt and Claudia Czado. Standardized drought indices: a novel univariate and multivariate approach. Applied Statistics, 67(3):643–664, April 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Brockhaus:2018:SRM

[379] Sarah Brockhaus, Andreas Fuest, Andreas Mayr, and Sonja Greven. Signal regression models for location, scale and shape with an application to stock returns. Applied Statistics, 67(3):665– 686, April 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Abrams:2018:MTV

[380] Steven Abrams, Andreas Wienke, and Niel Hens. Modelling time varying heterogeneity in recurrent infection processes: an application to serological data. Applied Statistics, 67(3):687– 704, April 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jones-Todd:2018:SMM

[381] Charlotte M. Jones-Todd, Ben Swallow, Janine B. Illian, and Mike Toms. A spatiotemporal multispecies model of a semicontinuous response. Applied Statistics, 67(3):705–722, April 2018. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chu:2018:BBL

[382] Yiyi Chu and Ying Yuan. BLAST:

Bayesian latent subgroup design for basket trials accounting for patient heterogeneity. Applied Statistics, 67(3):723– 740, April 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2018:IId

[383] Anonymous. Issue information. Applied Statistics, 67(4):741–742, August 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Laber:2018:OTA

[384] Eric B. Laber, Nick J. Meyer, Brian J.

Reich, Krishna Pacifici, Jaime A. Collazo, and John M. Drake. Optimal treatment allocations in space and time for on-line control of an emerging infectious disease. Applied Statistics, 67(4):743– 789, August 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Liu:2018:UAC

[385] Yang Liu, Matías Salibián-Barrera, Ruben H. Zamar, and James V. Zidek. Using artificial censoring to improve extreme tail quantile estimates. Applied Statistics, 67(4):791–812, August 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bonnet:2018:IHE

[386] Anna Bonnet, Céline Lévy-Leduc, Elisabeth Gassiat, Roberto Toro, and Thomas Bourgeron. Improving heritability estimation by a variable selection approach in sparse high dimensional linear mixed models. Applied Statistics, 67(4):813–839, August 2018. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

OKeeffe:2018:CMM

[387] Aidan G. O’Keeffe, Li Su, and Vernon T. Farewell. Correlated multistate models for multiple processes: an application to renal disease progression in systemic lupus erythematosus. Applied Statistics, 67(4):841–860, August 2018. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fabrizi:2018:BSA

[388] Enrico Fabrizi, Maria Rosaria Ferrante, and Carlo Trivisano. Bayesian small area estimation for skewed business survey variables. Applied Statistics, 67(4): 861–879, August 2018. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sim:2018:NPB

[389] Gyuseok Sim, Ho Kim, Antonella Zanobetti, Joel Schwartz, and Yeonseung Chung. Non-parametric Bayesian multivariate metaregression: an application in environmental epidemiology. Applied Statistics, 67(4):881–896, August 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bouveyron:2018:FLB

[390] Charles Bouveyron, Laurent Bozzi, Julien Jacques, and François-Xavier Jollois. The functional latent block model for the co-clustering of electricity consumption curves. Applied Statistics, 67 (4):897–915, August 2018. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ertefaie:2018:DTE

[391] Ashkan Ertefaie, Jesse Y. Hsu, Lindsay C. Page, and Dylan S. Small. Discovering treatment effect heterogeneity through post-treatment variables with application to the effect of class size on mathematics scores. Applied Statistics, 67(4):917–938, August 2018. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cattelan:2018:MLR

[392] Manuela Cattelan and Cristiano Varin.

Marginal logistic regression for spatially clustered binary data. Applied Statistics, 67(4):939–959, August 2018. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Yan:2018:FPC

[393] Fangrong Yan, Xiao Lin, Ruosha Li, and Xuelin Huang. Functional principal components analysis on moving time windows of longitudinal data: dynamic prediction of times to event. Applied Statistics, 67(4):961–978, August 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Donat:2018:SEP

[394] Francesco Donat and Giampiero Marra. Simultaneous equation penalized likelihood estimation of vehicle accident injury severity. Applied Statistics, 67(4): 979–1001, August 2018. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shand:2018:SVA

[395] Lyndsay Shand, Bo Li, Trevor Park, and Dolores Albarracín. Spatially varying auto-regressive models for prediction of new human immunodeficiency virus diagnoses. Applied Statistics, 67(4):1003– 1022, August 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Brown:2018:EMP

[396] J. L. Brown and L. B. Hund. Estimating material properties under extreme conditions by using Bayesian model calibration with functional outputs. Applied Statistics, 67(4):1023– 1045, August 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hong:2018:PCP

[397] Hwanhee Hong, Haoda Fu, and Bradley P. Carlin. Power and commensurate priors for synthesizing aggregate and individual patient level data in network meta-analysis. Applied Statistics, 67(4):1047–1069, August 2018. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Goldstein:2018:BMW

[398] Harvey Goldstein, James Carpenter, and Michael G. Kenward. Bayesian models for weighted data with missing values: a bootstrap approach. Applied Statistics, 67(4):1071–1081, August 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Barcella:2018:MCB

[399] William Barcella, Maria De Iorio, and James Malone-Lee. Modelling correlated binary variables: an application to lower urinary tract symptoms. Applied Statistics, 67(4):1083–1100, August 2018. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2018:IIe

[400] Anonymous. Issue information. Applied Statistics, 67(5):1101–1102, November 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pigoli:2018:SAA

[401] Davide Pigoli, Pantelis Z. Hadjipantelis, John S. Coleman, and John A. D. Aston. The statistical analysis of acoustic phonetic data: exploring differences between spoken Romance languages. Applied Statistics, 67(5):1103– 1145, November 2018. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Olsen:2018:SIM

[402] Niels Lundtorp Olsen, Bo Markussen, and Lars Lau Raket. Simultaneous inference for misaligned multivariate functional data. Applied Statistics, 67(5): 1147–1176, November 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Copas:2018:RSO

[403] John B. Copas, Dan Jackson, Ian R. White, and Richard D. Riley. The role of secondary outcomes in multivariate meta-analysis. Applied Statistics, 67 (5):1177–1205, November 2018. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lewis:2018:PRU

[404] Nicole H. Lewis, David B. Hitchcock, Ian L. Dryden, and John R. Rose. Peptide refinement by using a stochastic search. Applied Statistics, 67(5):1207– 1236, November 2018. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rajala:2018:DMI

[405] T. Rajala, D. J. Murrell, and S. C. Olhede. Detecting multivariate interactions in spatial point patterns with Gibbs models and variable selection. Applied Statistics, 67(5):1237– 1273, November 2018. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Verbelen:2018:UPP

[406] Roel Verbelen, Katrien Antonio, and Gerda Claeskens. Unravelling the predictive power of telematics data in car insurance pricing. Applied Statistics, 67 (5):1275–1304, November 2018. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Reinhart:2018:SEP

[407] Alex Reinhart and Joel Greenhouse. Self-exciting point processes with spatial covariates: modelling the dynamics of crime. Applied Statistics, 67(5):1305– 1329, November 2018. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ciarleglio:2018:CTD

[408] Adam Ciarleglio, Eva Petkova, Todd Ogden, and Thaddeus Tarpey. Constructing treatment decision rules based on scalar and functional predictors when moderators of treatment effect are unknown. Applied Statistics, 67(5):1331– 1356, November 2018. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bharath:2018:RIB

[409] Karthik Bharath, Sebastian Kurtek, Arvind Rao, and Veerabhadran Baladandayuthapani. Radiologic imagebased statistical shape analysis of brain tumours. Applied Statistics, 67(5):1357– 1378, November 2018. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bauer:2018:SST

[410] Cici Bauer and Jon Wakefield. Stratified space-time infectious disease modelling, with an application to hand, foot and mouth disease in China. Applied Statistics, 67(5):1379–1398, November 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cagnone:2018:MLV

[411] Silvia Cagnone and Cinzia Viroli. Multivariate latent variable transition models of longitudinal mixed data: an analysis on alcohol use disorder. Applied Statistics, 67(5):1399–1418, November 2018. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

vanWieringen:2018:TPA

[412] Wessel N. van Wieringen, Carel F. W. Peeters, Renee X. de Menezes, and Mark A. van de Wiel. Testing for pathway (in)activation by using Gaussian graphical models. Applied Statistics, 67 (5):1419–1436, November 2018. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zhang:2018:NLM

[413] Hongbin Zhang and Lang Wu. A nonlinear model for censored and mismeasured time varying covariates in survival models, with applications in human immunodeficiency virus and acquired immune deficiency syndrome studies. Applied Statistics, 67(5):1437– 1450, November 2018. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2018:IAV

[414] Anonymous. Index of authors, volume 67, 2018. Applied Statistics, 67(5):1451– 1456, November 2018. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2018:CV

[415] Anonymous. Contents of volume 67, 2018. Applied Statistics, 67(5):1457– 1459, November 2018. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2019:IIa

[416] Anonymous. Issue information. Applied Statistics, 68(1):1–2, January 2019. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pitkin:2019:DPM

[417] James Pitkin, Gordon Ross, and Ioanna Manolopoulou. Dirichlet process mixtures of order statistics with applications to retail analytics. Applied Statistics, 68 (1):3–28, January 2019. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hilton:2019:PUM

[418] Jason Hilton, Erengul Dodd, Jonathan J.

Forster, and Peter W. F. Smith. Projecting UK mortality by using Bayesian generalized additive models. Applied Statistics, 68(1):29–49, January 2019. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chang:2019:CMC

[419] Kai-Lan Chang and Serge Guillas. Computer model calibration with large nonstationary spatial outputs: application to the calibration of a climate model. Applied Statistics, 68(1):51–78, January 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chen:2019:DLI

[420] Yin-Hsiu Chen, Bhramar Mukherjee, and Veronica J. Berrocal. Distributed lag interaction models with two pollutants. Applied Statistics, 68(1):79– 97, January 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kim:2019:SCD

[421] Heeyoung Kim, Rong Duan, Sungil Kim, Jaehwan Lee, and Guang-Qin Ma. Spatial cluster detection in mobility networks: a copula approach. Applied Statistics, 68(1):99–120, January 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cheng:2019:IRP

[422] Wenting Cheng, Jeremy M. G. Taylor, Tian Gu, Scott A. Tomlins, and Bhramar Mukherjee. Informing a risk prediction model for binary outcomes with external coefficient information. Applied Statistics, 68(1):121–139, January 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Behrouzi:2019:DES

[423] Pariya Behrouzi and Ernst C. Wit. Detecting epistatic selection with partially observed genotype data by using copula graphical models. Applied Statistics, 68(1):141–160, January 2019. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Boys:2019:RTM

[424] Richard J. Boys and Peter M. Philipson. On the ranking of test match batsmen. Applied Statistics, 68(1):161– 179, January 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dinsdale:2019:MPS

[425] Daniel Dinsdale and Matias Salibian-Barrera. Methods for preferential sampling in geostatistics. Applied Statistics, 68(1):181–198, January 2019. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gomez-Rubio:2019:MPI

[426] Virgilio Gómez-Rubio and Francisco Palmí-Perales. Multivariate posterior inference for spatial models with the integrated nested Laplace approximation. Applied Statistics, 68(1):199– 215, January 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Samartsidis:2019:BLG

[427] Pantelis Samartsidis, Claudia R. Eickhoff, Simon B. Eickhoff, Tor D. Wager, Lisa Feldman Barrett, Shir Atzil, Timothy D. Johnson, and Thomas E. Nichols. Bayesian log-Gaussian Cox process regression: applications to metaanalysis of neuroimaging working memory studies. Applied Statistics, 68(1): 217–234, January 2019. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tighiouart:2019:TSD

[428] Mourad Tighiouart. Two-stage design for phase I–II cancer clinical trials using continuous dose combinations of cytotoxic agents. Applied Statistics, 68(1): 235–250, January 2019. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hu:2019:TTB

[429] Zonghui Hu. Temporal trends of biomarkers and between-biomarker associations. Applied Statistics, 68(1):251– 264, January 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2019:IIb

[430] Anonymous. Issue information. Applied Statistics, 68(2):265–266, February 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wages:2019:PTI

[431] Nolan A. Wages and Ying Yuan. Preface to the themed issue on ‘Early phase clinical trial design methodology’. Applied Statistics, 68(2):267–269, February 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Clertant:2019:SDF

[432] M. Clertant and J. O’Quigley. Semiparametric dose finding methods: special cases. Applied Statistics, 68(2):271– 288, February 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mu:2019:GUM

[433] Rongji Mu, Ying Yuan, Jin Xu, Sumithra J. Mandrekar, and Jun Yin. gBOIN: a unified model-assisted phase I trial design accounting for toxicity grades, and binary or continuous end points. Applied Statistics, 68(2):289– 308, February 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wheeler:2019:BMF

[434] Graham M. Wheeler, Michael J. Sweeting, and Adrian P. Mander. A Bayesian model-free approach to combination therapy phase I trials using censored time-to-toxicity data. Applied Statistics, 68(2):309–329, February 2019. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Muenz:2019:PIT

[435] Daniel G. Muenz, Jeremy M. G. Taylor, and Thomas M. Braun. Phase I–II trial design for biologic agents using conditional auto-regressive models for toxicity and efficacy. Applied Statistics, 68(2): 331–345, February 2019. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mozgunov:2019:ITP

[436] Pavel Mozgunov and Thomas Jaki. An information theoretic phase I–II design for molecularly targeted agents that does not require an assumption of monotonicity. Applied Statistics, 68(2):347– 367, February 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ezzalfani:2019:JMB

[437] Monia Ezzalfani, Tomasz Burzykowski, and Xavier Paoletti. Joint modelling of a binary and a continuous outcome measured at two cycles to determine the optimal dose. Applied Statistics, 68(2): 369–384, February 2019. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lyu:2019:ATA

[438] Jiaying Lyu, Yuan Ji, Naiqing Zhao, and Daniel V. T. Catenacci. AAA: triple adaptive Bayesian designs for the identification of optimal dose combinations in dual-agent dose finding trials. Applied Statistics, 68(2):385–410, February 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Guo:2019:UBB

[439] Beibei Guo, Yeonhee Park, and Suyu Liu. A utility-based Bayesian phase I– II design for immunotherapy trials with progression-free survival end point. Applied Statistics, 68(2):411–425, February 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ursino:2019:DFD

[440] Moreno Ursino, Ying Yuan, Corinne Alberti, Emmanuelle Comets, Geraldine Favrais, Tim Friede, Frederike Lentz, Nigel Stallard, and Sarah Zohar. A dose finding design for seizure reduction in neonates. Applied Statistics, 68(2):427– 444, February 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kakurai:2019:DIV

[441] Yasuyuki Kakurai, Shuhei Kaneko, Chikuma Hamada, and Akihiro Hirakawa. Dose individualization and variable selection by using the Bayesian lasso in early phase dose finding trials. Applied Statistics, 68(2):445–460, February 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lee:2019:ONK

[442] Juhee Lee, Peter F. Thall, and Katy Rezvani. Optimizing natural killer cell doses for heterogeneous cancer patients on the basis of multiple event times. Applied Statistics, 68(2):461–474, February 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Li:2019:TDP

[443] Yimei Li, Ming Wang, and Ying Kuen Cheung. Treatment and dose prioritization in early phase platform trials of targeted cancer therapies. Applied Statistics, 68(2):475–491, February 2019. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2019:IIc

[444] Anonymous. Issue information. Applied Statistics, 68(3):493–494, April 2019.

CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lim:2019:FCA

[445] Yaeji Lim, Hee-Seok Oh, and Ying Kuen Cheung. Functional clustering of accelerometer data via transformed input variables. Applied Statistics, 68(3):495– 520, April 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Teng:2019:BAF

[446] Ming Teng, Farouk S. Nathoo, and Timothy D. Johnson. Bayesian analysis of functional magnetic resonance imaging data with spatially varying autoregressive orders. Applied Statistics, 68 (3):521–541, April 2019. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sorbye:2019:CPS

[447] Sigrunn H. Sørbye, Janine B. Illian, Daniel P. Simpson, David Burslem, and Håvard Rue. Careful prior specification avoids incautious inference for log-Gaussian Cox point processes. Applied Statistics, 68(3):543–564, April 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Peluso:2019:DWG

[448] Alina Peluso, Veronica Vinciotti, and Keming Yu. Discrete Weibull generalized additive model: an application to count fertility data. Applied Statistics, 68(3):565–583, April 2019. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gervini:2019:EPD

[449] Daniel Gervini and Manoj Khanal. Exploring patterns of demand in bike sharing systems via replicated point process models. Applied Statistics, 68(3):585– 602, April 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anderson:2019:RLM

[450] Gordon Anderson, Alessio Farcomeni, Maria Grazia Pittau, and Roberto Zelli. Rectangular latent Markov models for time-specific clustering, with an analysis of the wellbeing of nations. Applied Statistics, 68(3):603–621, April 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Huang:2019:ODE

[451] Yuanzhi Huang, Steven G. Gilmour, Kalliopi Mylona, and Peter Goos. Optimal design of experiments for non-linear response surface models. Applied Statistics, 68(3):623–640, April 2019. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Price-Williams:2019:DWD

[452] Matthew Price-Williams, Nick Heard, and Patrick Rubin-Delanchy. Detecting weak dependence in computer network traffic patterns by using higher criticism. Applied Statistics, 68(3):641–655, April 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chen:2019:CNP

[453] Jack Kuang Tsung Chen, Richard L. Valliant, and Michael R. Elliott. Calibrating non-probability surveys to estimated control totals using LASSO, with an application to political polling. Applied Statistics, 68(3):657–681, April 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lin:2019:RFR

[454] Chien-Wei Lin, Serena G. Liao, Peng Liu, Mei-Ling Ting Lee, Yong Seok Park, and George C. Tseng. RNASeqDesign: a framework for ribonucleic acid sequencing genomewide power calculation and study design issues. Applied Statistics, 68(3):683–704, April 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zhou:2019:PBS

[455] Tianjian Zhou, Peter Müller, Subhajit Sengupta, and Yuan Ji. PairClone: a Bayesian subclone caller based on mutation pairs. Applied Statistics, 68(3): 705–725, April 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zheng:2019:RIG

[456] Chao Zheng, Davide Ferrari, Michael Zhang, and Paul Baird. Ranking the importance of genetic factors by variableselection confidence sets. Applied Statistics, 68(3):727–749, April 2019. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zafari:2019:TMM

[457] Babak Zafari and Tahir Ekin. Topic modelling for medical prescription fraud and abuse detection. Applied Statistics, 68(3):751–769, April 2019. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zhu:2019:LLT

[458] Yayuan Zhu, Liang Li, and Xuelin Huang. Landmark linear transformation model for dynamic prediction with application to a longitudinal cohort study of chronic disease. Applied Statistics, 68 (3):771–791, April 2019. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

ONeill:2019:NCR

[459] Donal O’Neill. A new competing risks decomposition: application to the effect of cutting unemployment benefit on unemployment durations. Applied Statistics, 68(3):793–807, April 2019. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Xu:2019:BNP

[460] Yanxun Xu, Peter F. Thall, William Hua, and Borje S. Andersson. Bayesian non-parametric survival regression for optimizing precision dosing of intravenous busulfan in allogeneic stem cell transplantation. Applied Statistics, 68 (3):809–828, April 2019. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2019:IId

[461] Anonymous. Issue information. Applied Statistics, 68(4):829–830, August 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jalbert:2019:MER

[462] Jonathan Jalbert, Orla A. Murphy, Christian Genest, and Johanna G. Neslehová. Modelling extreme rain accumulation with an application to the 2011 Lake Champlain flood. Applied Statistics, 68(4):831–858, August 2019. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Davies:2019:IIA

[463] Vinny Davies, William T. Harvey, Richard Reeve, and Dirk Husmeier. Improving the identification of antigenic sites in the H1N1 influenza virus through accounting for the experimental structure in a sparse hierarchical Bayesian model. Applied Statistics, 68(4):859– 885, August 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Picchini:2019:BIS

[464] Umberto Picchini and Julie Lyng Forman. Bayesian inference for stochastic differential equation mixed effects models of a tumour xenography study. Applied Statistics, 68(4):887–913, August 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Flandre:2019:CKM

[465] Philippe Flandre and John O’Quigley.

Comparing Kaplan–Meier curves with delayed treatment effects: applications in immunotherapy trials. Applied Statistics, 68(4):915–939, August 2019. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Aflakparast:2019:DIB

[466] Mehran Aflakparast and Mathisca de Gunst. Data integrative Bayesian inference for mixtures of regression models. Applied Statistics, 68(4):941– 962, August 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chen:2019:MSI

[467] Cathy W. S. Chen, Khemmanant Khamthong, and Sangyeol Lee. Markov switching integer-valued generalized auto-regressive conditional heteroscedastic models for dengue counts. Applied Statistics, 68(4):963–983, August 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Martinez-Camblor:2019:IVP

[468] Pablo Martínez-Camblor, Todd A. MacKenzie, Douglas O. Staiger, Phillip P. Goodney, and A. James O’Malley. An instrumental variable procedure for estimating Cox models with nonproportional hazards in the presence of unmeasured confounding. Applied Statistics, 68(4):985–1005, August 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Markert:2019:DAF

[469] Karla Markert, Karolin Krehl, Carsten Gottschlich, and Stephan Huckemann. Detecting anisotropy in fingerprint growth. Applied Statistics, 68(4):1007– 1027, August 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Scheike:2019:MVC

[470] Thomas H. Scheike, Frank Eriksson, and Siri Tribler. The mean, variance and correlation for bivariate recurrent event data with a terminal event. Applied Statistics, 68(4):1029– 1049, August 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sestelo:2019:TCP

[471] Marta Sestelo and Javier Roca-Pardiñas. Testing critical points of nonparametric regression curves: application to the management of stalked barnacles. Applied Statistics, 68(4):1051– 1070, August 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Geraci:2019:AQR

[472] Marco Geraci. Additive quantile regression for clustered data with an application to children’s physical activity. Applied Statistics, 68(4):1071– 1089, August 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kunihama:2019:NPB

[473] Tsuyoshi Kunihama, Carolyn T.

Halpern, and Amy H. Herring. Nonparametric Bayes models for mixed scale longitudinal surveys. Applied Statistics, 68(4):1091–1109, August 2019. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lam:2019:NPO

[474] Chi Kin Lam, Ruitao Lin, and Guosheng Yin. Non-parametric overdose control for dose finding in drug combination trials. Applied Statistics, 68(4):1111– 1130, August 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dey:2019:EBG

[475] Rajib Dey, Noel Cadigan, and Nan Zheng. Estimation of the Von Bertalanffy growth model when ages are measured with error. Applied Statistics, 68(4):1131–1147, August 2019. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Watson:2019:HES

[476] Samuel I. Watson, Jo Sartori, Olalekan Uthman, and Richard J. Lilford. Health effects of sanitation facilities: a Bayesian semiparametric analysis of compositional data. Applied Statistics, 68(4): 1149–1166, August 2019. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lee:2019:JMC

[477] Youjin Lee, Mei-Cheng Wang, Katherine L. Grantz, and Rajeshwari Sundaram. Joint modelling of competing risks and current status data: an application to a spontaneous labour study. Applied Statistics, 68(4):1167– 1182, August 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tosetti:2019:CEC

[478] Elisa Tosetti and Veronica Vinciotti. A computationally efficient correlated mixed probit model for credit risk inference. Applied Statistics, 68(4):1183– 1204, August 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2019:IIe

[479] Anonymous. Issue information. Applied Statistics, 68(5):1205–1206, November 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Price:2019:ACI

[480] Michael J. Price, Cindy L. Yu, David A. Hennessy, and Xiaodong Du. Are actuarial crop insurance rates fair?: an analysis using a penalized bivariate B-spline method. Applied Statistics, 68(5):1207– 1232, November 2019. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fauvernier:2019:MDP

[481] Mathieu Fauvernier, Laurent Roche, Zoé Uhry, Laure Tron, Nadine Bossard, Laurent Remontet, and The Challenges in the Estimation of Net Survival Working Survival Group. Multi-dimensional penalized hazard model with continuous covariates: applications for studying trends and social inequalities in cancer survival. Applied Statistics, 68(5):1233– 1257, November 2019. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sansom:2019:SSM

[482] Philip G. Sansom, Daniel B. Williamson, and David B. Stephenson. State space models for non-stationary intermittently coupled systems: an application to the North Atlantic oscillation. Applied Statistics, 68(5):1259–1280, November 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cardot:2019:EFM

[483] Hervé Cardot, Guillaume Lecuelle, Pascal Schlich, and Michel Visalli. Estimating finite mixtures of semi-Markov chains: an application to the segmentation of temporal sensory data. Applied Statistics, 68(5):1281–1303, November 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mohammed:2019:BVS

[484] Shariq Mohammed, Dipak K. Dey, and Yuping Zhang. Bayesian variable selection using spike-and-slab priors with application to high dimensional electroencephalography data by local modelling. Applied Statistics, 68(5):1305– 1326, November 2019. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Selosse:2019:AQL

[485] Margot Selosse, Julien Jacques, Christophe Biernacki, and Florence Cousson-Gélie. Analysing a quality-of-life survey by using a coclustering model for ordinal data and some dynamic implications. Applied Statistics, 68(5):1327–1349, November 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kjaergaard:2019:FCD

[486] Søren Kjærgaard, Yunus Emre Ergemen, Malene Kallestrup-Lamb, Jim Oeppen, and Rune Lindahl-Jacobsen. Forecasting causes of death by using compositional data analysis: the case of cancer deaths. Applied Statistics, 68 (5):1351–1370, November 2019. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dai:2019:COT

[487] James Y. Dai and Michael LeBlanc. Case-only trees and random forests for exploring genotype-specific treatment effects in randomized clinical trials with dichotomous end points. Applied Statistics, 68(5):1371–1391, November 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Domenicano:2019:BUD

[488] I. Domenicano, S. Ventz, M. Cellamare, R. H. Mak, and L. Trippa. Bayesian uncertainty-directed dose finding designs. Applied Statistics, 68(5):1393– 1410, November 2019. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ott:2019:RBR

[489] Miles Q. Ott, Krista J. Gile, Matthew T.

Harrison, Lisa G. Johnston, and Joseph W. Hogan. Reduced bias for respondent-driven sampling: accounting for non-uniform edge sampling probabilities in people who inject drugs in Mauritius. Applied Statistics, 68(5):1411– 1429, November 2019. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sachs:2019:EPT

[490] Michael C. Sachs, Andrea Discacciati, Åsa H. Everhov, Ola Olén, and Erin E. Gabriel. Ensemble prediction of timeto-event outcomes with competing risks: a case-study of surgical complications in Crohn’s disease. Applied Statistics, 68(5):1431–1446, November 2019. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wan:2019:SAT

[491] Fang Wan, Andrew C. Titman, and Thomas F. Jaki. Subgroup analysis of treatment effects for misclassified biomarkers with time-to-event data. Applied Statistics, 68(5):1447– 1463, November 2019. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Andridge:2019:INI

[492] Rebecca R. Andridge, Brady T. West, Roderick J. A. Little, Philip S. Boonstra, and Fernanda Alvarado-Leiton. Indices of non-ignorable selection bias for proportions estimated from non-probability samples. Applied Statistics, 68(5):1465– 1483, November 2019. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tu:2019:PHD

[493] Chun-Chen Tu, Florence Forbes, Benjamin Lemasson, and Naisyin Wang. Prediction with high dimensional regression via hierarchically structured Gaussian mixtures and latent variables. Applied Statistics, 68(5):1485– 1507, November 2019. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Li:2019:TVA

[494] Xiaoye Li and Zhibiao Zhao. A time varying approach to the stock returninflation puzzle. Applied Statistics, 68 (5):1509–1528, November 2019. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Economou:2019:HSM

[495] Theodoros Economou and Matthew B. Menary. A hidden semi-Markov model for characterizing regime shifts in ocean density variability. Applied Statistics, 68(5):1529–1553, November 2019. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Davies:2019:FPI

[496] Vinny Davies, Umberto Noè, Alan Lazarus, Hao Gao, Benn Macdonald, Colin Berry, Xiaoyu Luo, and Dirk Husmeier. Fast parameter inference in a biomechanical model of the left ventricle by using statistical emulation. Applied Statistics, 68(5):1555–1576, November 2019. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Maity:2019:ISB

[497] Arnab Kumar Maity, Raymond J. Carroll, and Bani K. Mallick. Integration of survival and binary data for variable selection and prediction: a Bayesian approach. Applied Statistics, 68(5):1577– 1595, November 2019. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2019:IAV

[498] Anonymous. Index of authors, volume 68, 2019. Applied Statistics, 68(5):1597– 1603, November 2019. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:2019:CV

[499] Anonymous. Contents of volume 68, 2019. Applied Statistics, 68(5):1605– 1608, November 2019. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).